### v1

# StPR: Spatiotemporal Preservation and Routing for Exemplar-Free Video Class-Incremental Learning

Abstract
Video Class-Incremental Learning (VCIL) seeks to develop models that continuously learn new action categories over time without forgetting previously acquired knowledge. Unlike traditional Class-Incremental Learning (CIL), VCIL introduces the added complexity of spatiotemporal structures, making it particularly challenging to mitigate catastrophic forgetting while effectively capturing both frame-shared semantics and temporal dynamics. Existing approaches either rely on exemplar rehearsal, raising concerns over memory and privacy, or adapt static image-based methods that neglect temporal modeling. To address these limitations, we propose Spatiotemporal Preservation and Routing (StPR), a unified and exemplar-free VCIL framework that explicitly disentangles and preserves spatiotemporal information. First, we introduce Frame-Shared Semantics Distillation (FSSD), which identifies semantically stable and meaningful channels by jointly considering semantic sensitivity and classification contribution. These important semantic channels are selectively regularized to maintain prior knowledge while allowing for adaptation. Second, we design a Temporal Decomposition-based Mixture-of-Experts (TD-MoE), which dynamically routes task-specific experts based on their temporal dynamics, enabling inference without task ID or stored exemplars. Together, StPR effectively leverages spatial semantics and temporal dynamics, achieving a unified, exemplar-free VCIL framework. Extensive experiments on UCF101, HMDB51, and Kinetics400 show that our method outperforms existing baselines while offering improved interpretability and efficiency in VCIL. Code is available in the supplementary materials.
1 Introduction
Class-Incremental Learning (CIL) li2017learning ; belouadah2021comprehensive ; de2021continual ; masana2022class ; zhang2024center develops models that learn from a sequence of tasks without forgetting previous knowledge, recognizing an ever-growing set of classes without past task data or identifiers. A key challenge is catastrophic forgetting CatastrophicInterference1989McCloskey ; ConnectionistModels1990Ratcliff , where new knowledge overwrites old. While well studied for images, extending CIL to videos—Video Class-Incremental Learning (VCIL) park2021class ; villa2022vclimb —remains underexplored. VCIL differs from CIL by requiring continual learning of new categories while modeling frame-shared semantics and temporal dependencies, unlike CIL’s focus on static images. This spatiotemporal complexity is critical for understanding actions, motion, and scene dynamics in real-world applications like surveillance, driver monitoring, robotics, and lifelogging. Further, memory and privacy constraints often prohibit storing past video data, demanding continual learning without rehearsal.
Figure 1: (a) Prior methods rely on exemplar rehearsal or naively stack video and CIL modules. (b) Our StPR framework explicitly decouples and reuses spatiotemporal semantics to mitigate forgetting.
The central challenge of VCIL lies in mitigating catastrophic forgetting while effectively leveraging frame-shared semantics and temporal dynamics to incrementally learn new categories. Existing methods can be broadly categorized into two types, as illustrated in Figure 1(a): 1) Exemplar-based methods rebuffi2017icarl ; hou2019learning ; douillard2020podnet ; park2021class ; pei2022learning ; villa2022vclimb ; alssum2023just ; liang2024hypercorrelation store a portion of previous data (video clips, frames, or features) and apply rehearsal to reduce forgetting. However, storing exemplars incurs memory and privacy costs and typically emphasizes frame-level learning without explicitly modeling temporal dynamics. 2) CIL-based methods li2017learning ; dhar2019learning ; cheng2024stsp adapt algorithms developed for static images, using techniques like regularization or subspace projection. While avoiding exemplar storage, they often overlook spatiotemporal properties by flattening or underutilizing temporal features. In contrast, our method StPR (Figure 1(b)) explicitly decouples video features into Frame shared semantics and temporal dynamics, and reuses these decomposed components to enhance the model’s ability to adapt continually, thereby reducing forgetting without storing extensive exemplars.
Specifically, we propose a unified, exemplar-free VCIL framework that fully exploits the spatiotemporal nature of videos. Our method integrates both spatial semantic consistency and temporal variation to mitigate forgetting and improve adaptation across tasks. Separately, we introduce: 1) Frame-Shared Semantics Distillation (FSSD). To preserve frame-shared semantics and alleviate forgetting, we quantify the semantic importance of each channel using a combination of semantic sensitivity and classification contribution. This ensures that semantically meaningful and stable channels are preserved, achieving a better trade-off between stability and plasticity. 2) Temporal-Decomposition-based Mixture-of-Experts (TD-MoE). To exploit temporal dynamics for continual adaptation, we decouple task-specific temporal cues for each expert. At inference, expert routing depends solely on the temporal dynamics of the input, without requiring task identities or stored exemplars. This enables dynamic assignment of weights to experts according to the temporal dynamics of the input, facilitating incremental learning of new categories.
Our framework uniquely bridges the gap between video-specific spatiotemporal representation and class-incremental adaptation. By disentangling and leveraging both spatial semantic channel consistency and temporal dynamics, it offers an effective and explainable solution for continual video understanding. Our main contributions are: 1) We propose a Frame-Shared Semantics Distillation method (FSSD) that preserves frame-shared, semantically aligned spatial channels through semantic importance-aware regularization, optimizing the stability-plasticity trade-off in continual learningl; 2) We design a Temporal Decomposition based Mixture-of-Experts strategy (TD-MoE) that decomposes spatiotemporal features and uses temporal dynamics for expert combination, enabling task-id-free and dynamic adaptation; 3) We present a unified, exemplar-free VCIL framework that achieves state-of-the-art results on UCF101, HMDB51, and Kinetics400, demonstrating the effectiveness of integrating spatial semantics and temporal dynamics in VCIL.
2 Related Work
2.1 Class-Incremental Learning
Class-Incremental Learning (CIL) aims to enable models to continually learn new classes without forgetting previously learned ones. Existing approaches typically fall into three categories: (1) regularization-based methods kirkpatrick2017overcoming ; zenke2017continual ; xiang2022coarse ; zhou2023hierarchical , which constrain parameter updates to preserve prior knowledge, often via knowledge distillation li2017learning ; hou2019learning ; (2) replay-based methods bang2021rainbow ; chaudhry2018riemannian ; rebuffi2017icarl , which store or generate past data to reduce forgetting; and (3) structure-based methods serra2018overcoming ; mallya2018packnet ; mallya2018piggyback ; liang2024inflora ; yu2024boosting , which expand model capacity or isolate task-specific components. Recently, CIL combined with pre-trained vision transformers (ViTs) ermis2022memory ; CODAPromptCOntinual2023Smith ; wang2022dualprompt ; wang2022learning has achieved impressive results by leveraging transferable representations and modularity. Some methods fully fine-tune pre-trained backbones boschini2022transfer ; zhang2023slca , but this is computationally expensive. To address efficiency, parameter-efficient fine-tuning (PEFT) methods have been introduced. Prompt pool-based approaches wang2022learning ; CODAPromptCOntinual2023Smith ; wang2024hierarchical ; zhang2023slca maintain task-specific prompts, while adapter-based methods zhou2023revisiting ; tan2024semantically ; gao2024beyond ; liang2024inflora ; ExpandableSubspace2024Zhou adapt ViTs to new classes with minimal updates. While effective, most CIL strategies were originally developed for static image domains and do not generalize well to video-based scenarios, where temporal dynamics play a critical role.
2.2 Video Class-Incremental Learning
Action recognition has been widely explored with 2D CNNs using temporal aggregation lin2019tsm ; wang2016temporal and 3D CNNs for joint spatiotemporal modeling carreira2017quo ; tran2015learning . More recent work focuses on improving temporal sensitivity and efficiency feichtenhofer2020x3d ; fan2020rubiksnet . However, these models are trained in static setups and do not address continual adaptation or forgetting.
To address these challenges, Video Class-Incremental Learning (VCIL) extends conventional Class-Incremental Learning (CIL) to spatiotemporal data, introducing additional challenges such as managing temporal variations across tasks. Several recent methods, including TCD park2021class , FrameMaker pei2022learning , and HCE liang2024hypercorrelation , address this setting by storing videos or compressed exemplars. However, these strategies raise concerns related to memory efficiency and data privacy. While SMILE alssum2023just effectively extracts image features from individual frames, it does not explicitly capture temporal information, which may limit its ability to leverage the distinctive decision cues present in video data. Exemplar-free methods such as STSP cheng2024stsp mitigate forgetting via orthogonal subspace projections, but they largely adapt image-domain strategies to video tasks without fully exploiting spatiotemporal structures. In contrast, our approach explicitly decouples and models the spatiotemporal structure of videos. We propose a unified VCIL framework that (i) preserves spatial semantic consistency through Frame-Shared Semantics Distillation (FSSD), enhancing knowledge retention without requiring exemplars, and (ii) leverages temporal dynamics for expert routing via Temporal Decomposition-based Mixture-of-Experts (TD-MoE).
Figure 2: Overview of our proposed framework. (I) During training, the (a) Frame-Shared Semantics Distillation (FSSD) module retains past knowledge via frame-shared semantics importance-aware regularization. (II) At inference, the (b) Temporal Decomposition-based Mixture-of-Experts (TD-MoE) dynamically routes input videos to expert branches by leveraging the disentangled temporal component of the spatiotemporal representation, enabling adaptive predictions.
3 Method
Problem Definition: In the Video Class-Incremental Learning (VCIL) setting, a model is trained across $B$ stages with sequentially arriving datasets $\{\mathcal{D}^{1},\dots,\mathcal{D}^{B}\}$. Each dataset $\mathcal{D}^{b}=\{(V_{j}^{b},y_{j}^{b})\}_{j=1}^{|\mathcal{D}^{b}|}$ corresponds to the $b$-th task, where $V_{j}^{b}$ is the $j$-th video and $y_{j}^{b}$ is its class label. Here, videos primarily represent human action recognition scenarios, where the spatiotemporal dynamics capture motion patterns, subject interactions, and scene context. $|\mathcal{D}^{b}|$ represents the number of samples in the $b$-th task. Let $\mathcal{Y}^{b}$ be the label space of the $b$-th dataset. For all $b\neq b^{\prime}$, the label spaces are disjoint: $\mathcal{Y}^{t}\cap\mathcal{Y}^{b^{\prime}}=\varnothing$. The objective of VCIL is to incrementally train a model over $B$ tasks while maintaining high performance across all accumulated classes $\{\mathcal{Y}^{1},\mathcal{Y}^{2},\dots,\mathcal{Y}^{B}\}$.
Overall framework. We propose a unified, exemplar-free framework for Video Class-Incremental Learning (VCIL) built upon the CLIP model radford2021learning . Our goal is to mitigate catastrophic forgetting while effectively leveraging frame-shared semantics and temporal dynamics to incrementally learn new categories. The frozen visual encoder $\mathcal{F}(\cdot)$ extracts spatial features, while adapters $\mathcal{A}^{b}$ are updated for each task $b$. A spatiotemporal encoder $\mathcal{G}(\cdot)$ models temporal dynamics. Our framework introduces two key components: 1) Frame-Shared Semantics Distillation (FSSD) identifies semantically stable channels across frames by combining Semantic Sensitivity and Classification Score, applying selective regularization to preserve critical spatial semantics while maintaining plasticity. 2) Temporal Decomposition based Mixture-of-Experts (TD-MoE). To exploit temporal dynamics for continual adaptation, we decouple shared static components and temporal dynamics. During inference, temporal dynamics are used to assign dynamic weights to expert temporal encoders, enabling task-id-free adaptation without requiring task identifiers or stored exemplars.
3.1 Spatial and Spatiotemporal Encoder
Spatial Encoder.
The shared adapter module chen2022adaptformer $\mathcal{A}^{b}=\{\mathcal{A}_{l}^{b}\}_{l=1}^{N}$ is utilized with a frozen CLIP-ViT model with $N$ layers of transformer module, serving as the spatial extractor. An adapter is an encoder-decoder architecture embedded into the residual of each transformer layer, facilitating transfer learning and enhancing downstream task performance. Typically, it consists of a down-sampling MLP $\mathbf{W}_{down}\in\mathbb{R}^{d\times d_{h}}$, a ReLU activation $\bm{\phi}(\cdot)$, and an up-sampling MLP $\mathbf{W}_{up}\in\mathbb{R}^{d_{h}\times d}$, where $d$ is the input and output dimension, and $d_{h}$ is the hidden dimension. For adapter, if input is $\mathbf{v}_{i}^{\prime}\in\mathbb{R}^{d}$, which is the $i$-th frame-level feature after the Multi-Head Self-Attention and residual connection in CLIP ViT, the output of adapter is:
$\mathbf{v}^{a}_{i}=\bm{\phi}(\mathbf{v}_{i}^{\prime}\mathbf{W}_{down})\mathbf{W}_{up}.$ (1)
The spatial feature of $i$-th frame $\mathbf{V}_{i}^{s}=\mathcal{F}(V_{i};\mathcal{A}^{b}),\mathbf{V}_{i}^{s}\in\mathbb{R}^{d_{vt}}$, where $V_{i}$ is the $i$-th frame of the video, $d_{vt}$ is the dimension of the aligned video-text features.
Spatiotemporal Encoder. To obtain the spatiotemporal representation, we feed both the frame-level spatial features and a learnable [CLS] token into a multi-head self-attention based spatiotemporal encoder $\mathcal{G}(\cdot)$. Specifically, the input to $\mathcal{G}$ consists of the frame features $\mathbf{V}^{s}_{1},\dots,\mathbf{V}^{s}_{N_{f}}$ and a [CLS] token $\mathbf{V}^{s}_{cls}\in\mathbb{R}^{d_{vt}}$, where $N_{f}$ represents the number of sampled frames. The temporal encoder outputs the spatiotemporal feature $\mathbf{V}^{st}$, corresponding to the transformed [CLS] token:
$\mathbf{V}^{st}=\mathcal{G}([\mathbf{V}^{s}_{cls};\mathbf{V}^{s}_{1};\dots;\mathbf{V}^{s}_{N_{f}}])[0]\in\mathbb{R}^{d_{vt}},$ (2)
where $[0]$ selects the output associated with the [CLS] token after attention-based aggregation.
More detaiails for Spatial and Spatiotemporal Encoder are provided in appendix B.2 and B.3.
3.2 Frame-Shared Semantics Distillation
In VCIL, shared adapter modules inevitably drift in feature space when adapting to new tasks, leading to forgetting. Directly applying classic uniform-weighted distillation from CIL to video tasks ignores the differences in semantic importance and temporal stability across video features. This leads to a suboptimal balance between stability and plasticity. To address this, we propose Frame-Shared Semantics Distillation (FSSD), which leverages the unique property of videos: despite variations across frames, core semantics (e.g., action concepts) are consistently captured by a stable subset of channels across frames. By identifying and selectively regularizing these frame-shared informative channels, FSSD preserves prior knowledge while enabling flexible adaptation to new tasks.
Frame-Shared Semantics Distillation (FSSD). To mitigate semantic drift across tasks, we introduce a distillation loss weighted by the frame-shared semantics importance:
$\mathcal{L}_{FSSD}=\frac{1}{|\mathcal{D}^{b}|\cdot d_{vt}}\sum_{c}^{|\mathcal{C}_{b}|}\sum_{i}^{N_{c}}\sum_{j}^{d_{vt}}I_{b-1,c,j}\cdot\|\bar{V}_{b-1,c,i,j}^{s}-\bar{V}_{b,c,i,j}^{s}\|_{2}^{2},$ (3)
where $|\mathcal{C}_{b}|$ is the number of classes in the $b$-th task, and $N_{c}$ is the number of samples per class. $I_{b-1,c,j}$ denotes the frame-shared importance of the $j$-th channel for class $c$ from task $(b{-}1)$. $\bar{V}_{b-1,c,i,j}^{s}$ and $\bar{V}_{b,c,i,j}^{s}$ are the $j$-th channel outputs of the $i$-th sample in class $c$, extracted from the spatial encoders of task $(b{-}1)$ and $b$, respectively, calculated on current data. As the previous model is frozen during training, estimating semantic importance is essential for preserving prior knowledge across tasks.
Frame-Shared Semantics. To quantify the importance of frame-shared semantics, we assess each channel based on two criteria: 1). Semantic Sensitivity. It measures the responsiveness to activation changes, thereby reflecting its reliability in representing consistent semantic information. and 2) Classification Score. It reflects the channel’s contribution to the final classification.
For semantic sensitivity, we employ Fisher Information to estimate how sensitively a channel’s activation influences the output. As spatial features $\bar{\mathbf{V}}^{s}=\frac{1}{N_{f}}\sum_{i=1}^{N_{f}}\mathbf{V}^{s}_{i}$ aggregate frame-wise variations, the Central Limit Theorem suggests each channel’s distribution approximates a Gaussian (Belong to the same category). Thus, we assume the $j$-th channel activation for class $c$ follows (For simplicity, we omit the subscripts for task and sample.):
$\bar{V}^{s}_{c,j}\sim\mathcal{N}(\mu_{c,j},\sigma_{c,j}^{2}),$ (4)
where $\mu_{c,j}$ and $\sigma_{c,j}^{2}$ denote the mean and variance across frames. The Fisher Information $\mathcal{I}(\mu_{c,j})$ (Detailed derivations are provided in the appendix B.1) with respect to $\mu_{c,j}$ is computed as:
$\mathcal{I}(\mu_{c,j})=\mathbb{E}\left[\left(\frac{\bar{V}^{s}_{c,j}-\mu_{c,j}}{\sigma_{c,j}^{2}}\right)^{2}\right]=\frac{1}{\sigma_{j}^{4}}\mathbb{E}\left[(\bar{V}^{s}_{c,j}-\mu_{c,j})^{2}\right]=\frac{1}{\sigma_{j}^{4}}\cdot\sigma_{c,j}^{2}=\frac{1}{\sigma_{c,j}^{2}}.$ (5)
For classification score, we compute the cosine similarity between the spatial video feature $\bar{\mathbf{V}}^{s}_{c}\in\mathbb{R}^{d_{vt}}$ and its corresponding text feature $\mathbf{T}_{c}\in\mathbb{R}^{d_{vt}}$. Specifically, for the $j$-th channel, the classification score is defined as $\gamma_{c,j}=\frac{\bar{V}^{s}_{c,j}\cdot T_{c,j}}{\|\bar{\mathbf{V}}^{s}_{c}\|\cdot\|\mathbf{T}_{c}\|}$, where $T_{c,j}$ denotes the $j$-th dimension feature of $\mathbf{T}_{c}$. We then take the expectation of $\gamma_{c,j}$ across frames to obtain a stable channel-level contribution estimate:
$\mathbb{E}[\gamma_{c,j}]=\mathbb{E}\left[\frac{\bar{V}^{s}_{c,j}\cdot T_{c,j}}{\|\bar{\mathbf{V}}^{s}_{c}\|\cdot\|\mathbf{T}_{c}\|}\right]\propto\mathbb{E}\left[\frac{\bar{V}^{s}_{c,j}\cdot T_{c,j}}{\|\bar{\mathbf{V}}^{s}_{c}\|}\right]\approx\frac{T_{c,j}\cdot\mu_{c,j}}{\lambda},$ (6)
where $\|\bar{\mathbf{V}}^{s}_{c}\|\approx\lambda$ is treated as a constant after normalization.
Combining semantic sensitivity and classification score, the semantic importance for the $j$-th channel of the $c$-th class is defined as:
$I_{c,j}=\frac{T_{c,j}\cdot\mu_{c,j}}{\sigma_{c,j}^{2}}.$ (7)
After each task, we compute the frame-shared semantic importance for all classes and accumulate it as the distillation weight for the next task. FSSD ensures that, when learning new tasks, the model retains channels that are more semantically important for previous tasks while allowing less critical parameters to adapt to new knowledge, achieving a better balance between stability and plasticity.
3.3 Temporal Decomposition based Mixture-of-Experts
Given the high forgetting tendency of deep transformers and the unknown number of future tasks in VCIL, we allocate a dedicated spatiotemporal encoder for each task. As task ID are unavailable during inference, we introduce a routing mechanism to dynamically assign weights to the experts. This routing leverages the temporal patterns extracted from the video features, ensuring relevant experts contribute more to the final representation. This approach enables task ID-agnostic, exemplar-free inference, effectively leveraging the temporal structure of action videos to alleviate forgetting.
Task-Specific Expert. For each task, we train a dedicated expert based on the spatiotemporal encoder. The spatiotemporal features $\mathbf{V}^{st}\in\mathbb{R}^{d_{vt}}$ captured by each expert are computed as in Eq. 2. During inference, we introduce a dynamic routing mechanism, assigning weights to each expert based on temporal relevance, which allows us to obtain the Mixture-of-Experts (MoE) model feature.
Temporal Decomposition base-Router.
To design this routing mechanism based on temporal dynamics, we consider two aspects: 1) Temporal residuals. These reflect the subtle temporal differences within redundant frames. 2) Inter-frame information. This captures abstract temporal concepts between frames, based on the knowledge learned by each expert.
For temporal residuals, we observe that redundant frames, where backgrounds and subjects remain consistent, cause minimal variation between adjacent frames kim2024fusion ; liu2021no . This leads to short-term temporal stationarity, allowing frame features to be decomposed into a shared static component and small temporal residuals. We express each frame feature as $\mathbf{V}_{i}^{s}=\bar{\mathbf{v}}+\bm{\epsilon}_{i}$, where $\bar{\mathbf{v}}$ captures the shared static components (e.g., background, actor) and $\bm{\epsilon}_{i}$ represents zero-mean temporal fluctuations. The spatial representation is then the mean across frames:
$\bar{\mathbf{V}}^{s}=\bar{\mathbf{v}}+\bar{\bm{\epsilon}},\quad\bar{\bm{\epsilon}}=\frac{1}{N_{f}}\sum_{i=1}^{N_{f}}\bm{\epsilon}_{i}.$ (8)
For the inter-frame information, since the spatiotemporal feature $\mathbf{V}^{st}$ is computed by the attention module, it can be approximated as $\mathbf{V}^{st}\approx\sum_{i=1}^{N_{f}}a_{i}\cdot\mathbf{V}^{s}_{i}$, where $a_{i}$ is attention score. After normalization, we can obtain $\sum_{i=1}^{N_{f}}a_{i}=1$. Substituting this into the decomposition, we can simplify the expression in Eq. 8 as:
$\mathbf{V}^{st}=\bar{\mathbf{v}}+\sum_{i=1}^{N_{f}}a_{i}\cdot\bm{\epsilon}_{i}.$ (9)
Since $\bar{\mathbf{v}}$ is difficult to estimate, and to decouple the temporal residual $\bm{\epsilon}_{i}$ and inter-frame information $a_{i}$, we naturally address this by using the difference between $\mathbf{V}^{st}$ and $\bar{\mathbf{V}}^{s}$, effectively isolating the temporal dynamics, which can be represented as:
$\mathbf{V}^{tem}=\sum_{j=1}^{N_{f}}\left(a_{i}-\tfrac{1}{N_{f}}\right)\cdot\bm{\epsilon}_{i}.$ (10)
This formulation reveals that $\mathbf{V}^{tem}$ quantifies the deviation between the model’s attention-weighted temporal dynamics and the uniform temporal mean. It isolates temporal variations from static spatial semantics, allowing the model to use temporal cues for routing decisions while avoiding background interference, thus supporting our goal of mitigating forgetting and leveraging temporal dynamics for continual learning.
Inference. During inference, we first compute the decoupled temporal representation $\mathbf{V}^{tem}\in\mathbb{R}^{d_{vt}}$ for each input video. For all categories in the current task, we calculate the mean temporal representation and store it in the anchor pool as $\bar{\mathbf{V}}_{c}^{tem}\in\mathbb{R}^{d_{vt}}$, where $c$ represents the $c$-th class. For each expert $k$, we compute a similarity-based score as the router:
$r_{k}=\max_{c\in\mathcal{C}_{k}}\cos\left(\mathbf{V}_{k}^{tem},\bar{\mathbf{V}}_{c}^{tem}\right),$ (11)
where $\mathcal{C}_{k}$ represents the set of classes assigned to expert $k$. Then, we combine the adapter-tuned spatial features $\bar{\mathbf{V}}^{s}$ with the expert outputs weighted by $r_{k}$ as the final video representation:
$\mathbf{V}=\bar{\mathbf{V}}^{s}+\sum_{k}r_{k}\cdot\mathbf{V}_{k}^{cls}.$ (12)
The final video representation is matched with text embedding via cosine similarity for classification.
3.4 Loss Function and Optimization
The overall loss function in our framework includes: 1) a contrastive loss between the final video feature and text description features, serving as the overall classification loss; 2) the contrast loss between the video feature of clip output under adapter fine-tuning and the text feature as the spatial level optimization; and 3) the Frame-Shared Semantics Distillation loss, which mitigates forgetting in the shared adapter modules.
Contrastive Loss Formulation. We employ a symmetric contrastive loss for both video-to-text and text-to-video alignment. Given a batch of $N$ samples, let $\mathbf{V}_{i}$ and $\mathbf{T}_{j}$ denote the video and text features, respectively. The similarity between video $i$ and text $j$ is computed as the cosine similarity $S_{i,j}=\cos(\mathbf{V}_{i},\mathbf{T}_{j})$, forming a similarity matrix $S\in\mathbb{R}^{N\times N}$. Let $\mathbf{M}\in\{0,1\}^{N\times N}$ be the label mask, where $M_{i,j}=1$ if $y_{i}=y_{j}$ and $M_{i,j}=0$ otherwise. Then, the Video-to-text contrastive loss can be caculated by:
$\mathcal{L}_{\text{v2t}}=-\frac{1}{N}\sum_{i=1}^{N}\log\left(\frac{\sum_{j=1}^{N}M_{i,j}\cdot\exp(S_{i,j})}{\sum_{j=1}^{N}\exp(S_{i,j})+\varepsilon}\right),$ (13)
where $\varepsilon$ is a small constant added to avoid division by zero. Text-to-video contrastive loss is:
$\mathcal{L}_{\text{t2v}}=-\frac{1}{N}\sum_{j=1}^{N}\log\left(\frac{\sum_{i=1}^{N}M_{i,j}\cdot\exp(S_{i,j})}{\sum_{i=1}^{N}\exp(S_{i,j})+\varepsilon}\right).$ (14)
Symmetric total contrastive loss is :
$\mathcal{L}_{\text{Cont}}=\frac{1}{2}(\mathcal{L}_{\text{v2t}}+\mathcal{L}_{\text{t2v}}).$ (15)
For $\mathcal{L}_{\text{Cont}}^{\text{St}}$, the embeddings are the spatiotemporal video feature $\mathbf{V}^{st}$ and corresponding text features. For $\mathcal{L}_{\text{Cont}}^{\text{S}}$, we use the CLIP adapter feature $\bar{\mathbf{V}}^{s}$ and corresponding text features.
Total Loss. The overall training loss is:
$\mathcal{L}=\mathcal{L}_{\text{Cont}}^{\text{St}}+\mathcal{L}_{\text{Cont}}^{\text{S}}+w\cdot\mathcal{L}_{\text{FSSD}},$ (16)
where $w$ is hyper-parameter. This design ensures that both spatial and spatiotemporal semantics are aligned with text supervision, while the FSSD term selectively preserves critical frame-shared semantics to mitigate forgetting.
4 Experiments
Table 1: Video class-incremental results on UCF101 and HMDB51, with best results in bold.
Method Exemplars Venue UCF101 HMDB51
$10\times 5$s $5\times 10$s $2\times 25$s $5\times 5$s $1\times 25$s
iCaRL rebuffi2017icarl CVPR’17 $65.34$ $64.51$ $58.73$ $40.09$ $33.77$
LwFMC li2017learning TPAMI’18 $42.14$ $25.59$ $11.68$ $26.82$ $16.49$
LwM dhar2019learning CVPR’19 $43.39$ $26.07$ $12.08$ $26.97$ $16.50$
UCIR hou2019learning CVPR’19 $74.09$ $70.50$ $64.00$ $46.53$ $37.15$
PODNet douillard2020podnet ECCV’20 $74.37$ $73.75$ $71.87$ $48.78$ $46.62$
TCD park2021class ICCV’21 $77.16$ $75.35$ $74.01$ $50.36$ $46.66$
FrameMaker pei2022learning NeurIPS’22 $78.64$ $78.14$ $77.49$ $51.12$ $47.37$
STSP cheng2024stsp ECCV’24 $81.15$ $82.84$ $79.25$ $56.99$ $49.19$
HCE liang2024hypercorrelation AAAI’24 $80.01$ $78.81$ $77.62$ $52.01$ $48.94$
[…]
Table 2: Video class-incremental results on Kinetics400, with best results in bold.
Method Exemplars Venue Kinetics400-10s Kinetics400-20s
Acc $\uparrow$ BWF $\downarrow$ Acc $\uparrow$ BWF $\downarrow$
vCLIMB+BiC villa2022vclimb CVPR’22 $27.90$ $51.96$ $23.06$ $58.97$
vCLIMB+iCaRL villa2022vclimb CVPR’22 $32.04$ $38.74$ $26.73$ $42.25$
SMILE+BiC alssum2023just CVPR’23 $52.24$ $\mathbf{6.25}$ $48.22$ $\mathbf{0.31}$
SMILE+iCaRL alssum2023just CVPR’23 $46.58$ $7.34$ $45.77$ $4.57$
[…]
4.1 Experimental Details
Dataset. We evaluate our method on four benchmark datasets under the Video Class-Incremental Learning (VCIL) setting: UCF101 soomro2012ucf101 , HMDB51kuehne2011hmdb and Kinetics400 carreira2017quo . UCF101 and HMDB51 are widely used for human action recognition, containing 101 and 51 action classes respectively. Kinetics400 offers large-scale diversity with 400 action classes across 240,000 videos. We conduct all experiments under an exemplar-free setting. For fair comparison, we adopt the TCD protocol park2021class for UCF101 and HMDB51, where the model is pretrained on 51 and 26 base classes, and the remaining ones are distributed into 5, 10, or 25 tasks. For Kinetics400, we follow the vCLIMB benchmark villa2022vclimb , with 10- or 20-task splits.
Evaluation Metrics. We adopt two widely-used metrics to evaluate performance in VCIL: 1). Final Accuracy (Acc), which measures the overall classification accuracy on all learned classes after the final task is completed; 2). Backward Forgetting (BWF)villa2022vclimb , which quantifies the average drop in performance on previously learned tasks, reflecting how well the model retains past knowledge.
Implementation Details. All experiments are conducted on a single NVIDIA RTX 3090 GPU. We adopt the CLIP ViT-B/16 model radford2021learning as the backbone, with all its parameters frozen during training. The spatial and spatiotemporal encoders are the only trainable components in our framework. For optimization, we employ Stochastic Gradient Descent (SGD) with an initial learning rate of 0.01 and a batch size of 40. Each task is trained for 60 epochs in the first incremental session and 30 epochs in each subsequent session. The weighting hyperparameter $w$ in Eq. 16 is set to $1\times 10^{4}$. The multi-head self-attention module within the spatiotemporal encoder consists of three transformer layers, each employing two attention heads. Video clips are sampled using the TSN strategy wang2018temporal , selecting eight frames per video uniformly across the temporal dimension.
4.2 Main Results
Table. 1 and 2 report results on UCF101, HMDB51, and Kinetics400, covering different action complexities and temporal dynamics. Based on their strategies to mitigate forgetting, existing methods are categorized into two groups: 1) Exemplar-based methods.rebuffi2017icarl ; hou2019learning ; douillard2020podnet ; park2021class ; pei2022learning ; villa2022vclimb ; alssum2023just ; liang2024hypercorrelation . They store video clips, frames, or compressed features and apply rehearsal to reduce forgetting. However, these methods face scalability and privacy challenges due to their reliance on stored exemplars. 2) CIL-based methodsli2017learning ; dhar2019learning ; cheng2024stsp . This group adapts techniques from image-based class-incremental learning, such as unified distillation and subspace projection, without storing exemplars. While avoiding exemplar storage, their performance tends to be lower, especially as task difficulty increases and lacking explainable spatiotemporal disentanglement. In contrast, Our method (StPR) without storing exemplars, surpasses all baselines across datasets and settings. On UCF101 and HMDB51, it achieves 85.79% and 63.04% in the most challenging protocols, and outperforms SMILE+BiC and STSP on Kinetics400. This demonstrates the effectiveness of our core design: mitigating catastrophic forgetting by explicitly disentangling and reusing frame-shared semantics (via FSSD) and temporal dynamics (via TD-MoE) to enable robust continual learning.
Table 3: Ablation Study on UCF101 and HMDB51, with best results in bold.
Idx $\mathcal{A}^{b}$ FSSD TD-MoE UCF101($5\times 10$s) HMDB51($5\times 5$s) UCF101($10\times 5$s) HMDB51($25\times 1$s)
Acc $\uparrow$ BWF $\downarrow$ Acc $\uparrow$ BWF $\downarrow$ Acc $\uparrow$ BWF $\downarrow$ Acc $\uparrow$ BWF $\downarrow$
1 70.62 7.01 43.30 $\mathbf{6.18}$ 70.14 $\mathbf{5.33}$ 43.71 8.74
2 77.55 5.76 53.23 7.73 77.63 4.86 54.63 9.01
3 79.33 12.38 56.12 11.75 85.94 7.40 62.54 10.20
4 83.07 10.52 57.47 21.30 88.03 8.39 64.71 21.72
5 $\mathbf{85.79}$ $\mathbf{5.63}$ $\mathbf{63.04}$ 11.04 $\mathbf{88.52}$ $6.39$ $\mathbf{69.61}$ $\mathbf{7.02}$
Figure 3: Task-wise ablation analysis across incremental tasks on UCF101 and HMDB51.
4.3 Ablation Study
We perform ablation studies to evaluate the contribution of each component: the adapter tuning ($\mathcal{A}^{b}$), Frame-Shared Semantics Distillation (FSSD), and Temporal Decomposition-based Mixture-of-Experts (TD-MoE). Results are summarized in Table 3 and visualized across tasks in Figure 3. The baseline (pretrained CLIP) model exhibits limited performance on downstream tasks, as it lacks adaptation to new task-specific categories. Introducing FSSD alone moderately improves performance by preserving spatial semantics and reducing semantic drift, while TD-MoE independently enhances adaptation by leveraging temporal dynamics. However, using either module alone yields suboptimal performance. Combining adapter tuning with TD-MoE provides further improvements but still lacks sufficient stability in preserving spatial semantics. The full model (StPR), integrating both FSSD and TD-MoE, achieves the most stable performance across tasks. As shown in Figure 3, it maintains highest accuracy over time and effectively mitigates forgetting, demonstrating the complementary strengths of spatial semantic preservation and temporal dynamic modeling.
4.4 Further analysis
Analysis of temporal-decomposition routing strategies.
Table 4: MoE Method Results on UCF101 and HMDB51, with best results in bold.
Method UCF101($5\times 10$s) HMDB51($5\times 5$s) UCF101($10\times 5$s) HMDB51($25\times 1$s)
Acc $\uparrow$ BWF $\downarrow$ Acc $\uparrow$ BWF $\downarrow$ Acc $\uparrow$ BWF $\downarrow$ Acc $\uparrow$ BWF $\downarrow$
Avg-MoE $81.14$ 9.95 $59.46$ 8.49 84.04 9.60 62.69 9.14
CLIP-MoE $83.59$ 7.85 $58.82$ 10.01 85.80 7.27 65.43 8.08
Adapter-MoE $83.26$ 6.07 $61.99$ $\mathbf{7.75}$ 84.31 7.82 65.57 7.68
TD-MoE(Ours) $\mathbf{85.79}$ $\mathbf{5.63}$ $\mathbf{63.04}$ 11.04 $\mathbf{88.52}$ $\mathbf{6.39}$ $\mathbf{69.61}$ $\mathbf{7.02}$
Table 5: Distillation Method Results on UCF101 and HMDB51, with best results in bold.
Method UCF101($5\times 10$s) HMDB51($5\times 5$s) UCF101($10\times 5$s) HMDB51($25\times 1$s)
Acc $\uparrow$ BWF $\downarrow$ Acc $\uparrow$ BWF $\downarrow$ Acc $\uparrow$ BWF $\downarrow$ Acc $\uparrow$ BWF $\downarrow$
w/o Distillation $83.07$ 10.52 57.47 21.30 88.03 8.39 64.71 21.72
Distillation $84.27$ 7.45 $61.74$ 13.33 88.12 7.95 67.54 13.38
FSSD(Ours) $\mathbf{85.79}$ $\mathbf{5.63}$ $\mathbf{63.04}$ $\mathbf{11.04}$ $\mathbf{88.52}$ $\mathbf{6.39}$ $\mathbf{69.61}$ $\mathbf{7.02}$
Table 4 compares our TD-MoE with several alternative Mixture-of-Experts (MoE) jacobs1991adaptive routing strategies. Simple averaging (Avg-MoE) and static weight assignments—including CLIP-MoE, which uses frozen CLIP visual features for routing, and Adapter-MoE, which uses adapter-tuned CLIP features—provide moderate improvements but fail to dynamically leverage task-specific temporal cues, often resulting in higher forgetting. In contrast, TD-MoE enables adaptive expert weighting based on temporal dynamics, consistently improving both accuracy and stability across tasks. This highlights the importance of modeling temporal variability explicitly, rather than relying on static or feature-agnostic routing.
Effectiveness of FSSD over Uniform Distillation. Table 5 compares our FSSD method with the no-distillation baseline (w/o Distillation) and standard uniform distillation (Distillation) across four VCIL settings. While uniform distillation improves accuracy and reduces backward forgetting (BWF) over the naive baseline, FSSD consistently outperforms both, achieving the highest accuracy and lowest BWF in all settings. These results highlight the benefit of selectively preserving frame-shared semantics, validating the importance-aware design of FSSD for continual video learning.
Table 6: Analysis of hyperparameter $w$ on HMDB51 and UCF101 datasets. Best results are in bold.
Dataset Metric $1\times 10^{3}$ $1\times 10^{4}$ $2.5\times 10^{4}$ $5\times 10^{4}$ $1\times 10^{5}$
HMDB($5\times 5$s) Acc $\uparrow$ 59.51 63.04 62.10 62.29 63.05
BWF $\downarrow$ 16.90 11.04 11.14 12.32 12.77
UCF101($5\times 10$s) Acc $\uparrow$ 85.10 85.79 84.54 84.88 84.16
BWF $\downarrow$ 7.93 5.63 6.95 5.67 6.83
Analysis of hyper-parameter. Table. 6 analyzes the sensitivity of the hyper-parameter $w$, which controls the strength of FSSD regularization. Across a wide range of values, our framework maintains stable performance, indicating robustness to hyper-parameter variations. Moderate $w$ values achieve the best trade-off between knowledge retention and adaptability, avoiding under-regularization or excessive constraint on model plasticity. For more experiments, see the appendix C.
5 Conclusion
In this work, we propose StPR, a unified and exemplar-free framework for Video Class-Incremental Learning (VCIL) to tackle the spatiotemporal challenges in continual video learning. By disentangling spatial semantics and temporal dynamics, StPR effectively mitigates catastrophic forgetting without relying on stored exemplars. Our method combines Frame-Shared Semantics Distillation (FSSD), which selectively preserves meaningful and stable semantic channels, protecting model’s plasticity. Temporal-Decomposition-based Mixture-of-Experts (TD-MoE), adaptively routes inputs based on temporal cues, reducing forgetting in deep networks. Extensive experiments on UCF101, HMDB51, and Kinetics400 validate the effectiveness and efficiency of our approach, establishing new state-of-the-art results for continual video recognition. In future work, we plan to explore more realistic application scenarios, such as open-world settings, and investigate the deployment of our method on resource-constrained edge devices.
Appendix
Appendix A Algorithm
A.1 Algorithm of FSSD
Input: Current task data $\mathcal{D}_{b}$; frozen model from task $b{-}1$; text features $\{\mathbf{T}_{c}\}$
Output: FSSD loss $\mathcal{L}_{\text{FSSD}}$
for each class $c\in\mathcal{Y}_{b}$ do
       Compute mean spatial features:
       $\bar{\mathbf{V}}^{s}_{b,c}=\frac{1}{N_{f}}\sum_{i=1}^{N_{f}}\mathbf{V}^{s}_{b,c,i}$ ;
        // Current model
       $\bar{\mathbf{V}}^{s}_{b-1,c}=\frac{1}{N_{f}}\sum_{i=1}^{N_{f}}\mathbf{V}^{s}_{b-1,c,i}$ ;
        // Previous model
      for each channel $j=1$ to $d$ do
             Estimate $\mu_{c,j},\sigma^{2}_{c,j}$ from $\mathbf{V}^{s}_{b-1,c}$ ;
              // Across frames
            Compute semantic sensitivity:
             $\mathcal{I}(\mu_{c,j})=\frac{1}{\sigma^{2}_{c,j}}$ ;
              // Fisher Information
            Compute classification contribution:
             $\mathbb{E}[\gamma_{c,j}]\propto\frac{T_{c,j}\cdot\mu_{c,j}}{\lambda}$ ;
              // Cosine-aligned score
            Compute importance score:
             $I_{b-1,c,j}=\frac{T_{c,j}\cdot\mu_{c,j}}{\sigma^{2}_{c,j}}$ ;
              // Weighted relevance
Compute weighted distillation loss:
$\mathcal{L}_{FSSD}=\frac{1}{|\mathcal{D}^{b}|\cdot d_{vt}}\sum_{c}^{|\mathcal{C}_{b}|}\sum_{i}^{N_{c}}\sum_{j}^{d_{vt}}I_{b-1,c,j}\cdot\|\bar{V}_{b-1,c,i,j}^{s}-\bar{V}_{b,c,i,j}^{s}\|_{2}^{2}$
return $\mathcal{L}_{\text{FSSD}}$
Algorithm 1 Frame-Shared Semantics Distillation (FSSD)
A.2 Algorithm of TD-MoE
Input: Video frames $\{\mathbf{x}^{\text{V}}_{i}\}_{i=1}^{N_{f}}$;
    Task-specific experts $\{\mathcal{G}_{k}\}_{k=1}^{K}$;
    Temporal anchors $\{\bar{\mathbf{V}}^{tem}_{c}\}_{c=1}^{C}$ for current task
Output: Final representation $\mathbf{V}$
Compute spatial mean: $\bar{\mathbf{V}}^{s}=\frac{1}{N_{f}}\sum_{i=1}^{N_{f}}\mathbf{x}^{\text{V}}_{i}$ ;
  // Mean of frame features
for each expert $k=1$ to $K$ do
       Concatenate CLS token and frame features
       $\mathbf{V}_{k}^{st}=\mathcal{G}_{k}([\mathbf{x}^{\text{V}}_{\text{CLS}};\mathbf{x}^{\text{V}}_{1};\dots;\mathbf{x}^{\text{V}}_{N_{f}}])[0]$ ;
        // Spatiotemporal feature
      Compute temporal representation:
       $\mathbf{V}_{k}^{tem}=\mathbf{V}_{k}^{st}-\bar{\mathbf{V}}^{s}$ ;
        // Temporal decomposition
      Compute routing score:
       $r_{k}=\max_{c\in\mathcal{Y}_{k}}\cos(\mathbf{V}_{k}^{tem},\bar{\mathbf{V}}^{tem}_{c})$ ;
        // Similarity to temporal anchors
Compute final representation:
$\mathbf{V}=\bar{\mathbf{V}}^{s}+\sum_{k=1}^{K}r_{k}\cdot\mathbf{V}_{k}^{st}$ ;
  // Residual fusion
return $\mathbf{V}$
Algorithm 2 Temporal Decomposition based Mixture-of-Experts Inference
Frame-Shared Semantics Distillation (FSSD).
Algorithm 1 mitigates forgetting by selectively preserving spatial feature channels that are semantically important and temporally stable across frames. Importance is computed per channel using two criteria: (1) Semantic sensitivity, quantified by Fisher Information, and (2) Classification contribution, measured by cosine similarity with text features. These weights are used in a weighted distillation loss between the frozen previous model and the current task model, enabling exemplar-free knowledge retention while allowing plasticity.
Temporal Decomposition-based Mixture-of-Experts (TD-MoE).
Algorithm 2 routes video inputs to task-specific experts based on temporal relevance. Each expert encodes spatiotemporal features from video frames, from which temporal dynamics are isolated via residual decomposition. Temporal features are then compared to precomputed class anchors to compute routing scores. Final representations are generated by combining the expert outputs with the spatial feature via residual fusion. This enables dynamic, task ID-agnostic inference driven by temporal structure.
Appendix B Theoretical Supplement
B.1 Frame-Shared Semantics
Semantic Sensitivity. The Fisher Information with respect to the mean parameter $\mu_{j}$ is defined as:
$\mathcal{I}_{j}(\mu_{c,j})=\mathbb{E}_{\bar{V}^{s}_{c,j}}\left[\left(\frac{\partial}{\partial\mu_{c,j}}\log p(\bar{V}^{s}_{c,j};\mu_{c,j})\right)^{2}\right],$ (17)
where $p(\bar{V}^{s}_{c,j};\mu_{c,j})$ is the probability density function of the Gaussian:
$p(\bar{V}^{s}_{c,j};\mu_{c,j})=\frac{1}{\sqrt{2\pi\sigma_{c,j}^{2}}}\exp\left(-\frac{(\bar{V}^{s}_{c,j}-\mu_{c,j})^{2}}{2\sigma_{c,j}^{2}}\right).$ (18)
Taking the derivative of the log-likelihood with respect to $\mu_{i}$:
$\log p(\bar{V}^{s}_{c,j};\mu_{c,j})=-\frac{1}{2}\log(2\pi\sigma_{c,j}^{2})-\frac{(\bar{V}^{s}_{c,j}-\mu_{c,j})^{2}}{2\sigma_{i}^{2}},$ (19)
$\frac{\partial}{\partial\mu_{c,j}}\log p(\bar{V}^{s}_{c,j};\mu_{c,j})=\frac{\bar{V}^{s}_{c,j}-\mu_{c,j}}{\sigma_{c,j}^{2}}.$ (20)
Then, the Fisher Information becomes:
$\mathcal{I}(\mu_{c,j})=\mathbb{E}\left[\left(\frac{\bar{V}^{s}_{c,j}-\mu_{c,j}}{\sigma_{c.j}^{2}}\right)^{2}\right]=\frac{1}{\sigma_{j}^{4}}\mathbb{E}\left[(\bar{V}^{s}_{c,j}-\mu_{c,j})^{2}\right]=\frac{1}{\sigma_{j}^{4}}\cdot\sigma_{c,j}^{2}=\frac{1}{\sigma_{c,j}^{2}}.$ (21)
Thus, we obtain:
$\mathcal{I}(\mu_{c,j})=\frac{1}{\sigma_{c,j}^{2}}.$ (22)
Classification Contribution. As we use cosine distance as classification basis, the $c$-th classification decision score $\gamma_{c,j}$ of the $j$-th channel is modeled as:
$\gamma_{c,j}=\frac{\bar{V}^{s}_{c,j}\cdot T_{c,j}}{\|\bar{\bm{V}}_{c}^{s}\|\cdot\|\bm{T}_{c}\|}\propto\frac{\bar{V}^{s}_{c,j}\cdot T_{c,j}}{\|\bar{\bm{V}}_{c}^{s}\|}.$ (23)
Therefore, the score function s can be approximately written as:
$\gamma_{c,j}=\sum_{j}\bar{V}^{s}_{c,j}\cdot\alpha_{j},\quad\alpha_{j}=\frac{T_{c,j}}{\|\bar{\bm{V}}_{c}^{s}\|}\approx\frac{T_{c,j}}{\lambda}.$ (24)
For the expected score $\mathbb{E}[\gamma_{c,j}]$, it can be calculated as:
$\mathbb{E}[\gamma_{c,j}]\propto\mathbb{E}\left[\frac{\sum_{i}\bar{V}^{s}_{c,j}\cdot T_{c,j}}{\|\bar{\mathbf{x}}_{v}\|}\right]=\sum_{j}{\alpha_{j}\cdot\mu_{i}}\approx\frac{T_{c,j}\cdot\mu_{c,j}}{\lambda}\cdot\frac{\mu_{c,j}}{\sigma_{c,j}^{2}}.$ (25)
Then, the joint measure of informativeness is:
$I_{c,j}\propto\alpha_{j}\cdot\mu_{c,j}\cdot\mathcal{I}(\mu_{c,j})\approx\frac{T_{c,j}}{\lambda}\cdot\frac{\mu_{c,j}}{\sigma_{c,j}^{2}}.$ (26)
This expression provides a theoretically principled and interpretable metric for frame-shared semantics. It reflects the intuition that an informative channel should (i) be strongly activated on average ($\mu_{c,j}$ large), and (ii) exhibit consistent activation patterns across samples ($\sigma_{c,j}^{2}$ small). Therefore, we define the importance of channel $j$ as:
$I_{c,j}=\frac{T_{c,j}\cdot\mu_{c,j}}{\sigma_{c,j}^{2}}.$ (27)
This formulation also aligns with the the signal-to-noise ratio theory (SNR), providing a unified theoretical justification.
B.2 Spatial Encoder
We adopt a frozen CLIP-ViT model enhanced with shared adapters as the spatial encoder to extract frame-level features. Let a video $V=\{V_{i}\}_{i=1}^{N_{f}}$ consist of $N_{f}$ uniformly sampled frames, where each frame $V_{i}$ is processed independently. The spatial encoder $\mathcal{F}(V_{i};\mathcal{A}^{b})$ contains $L$ transformer layers, each equipped with an adapter $\mathcal{A}_{\ell}^{b}$ inserted after the multi-head self-attention (MHSA) residual. The output of the encoder is a set of spatial features $\{\mathbf{v}_{i}^{s}\}_{i=1}^{N_{f}}$.
Transformer Block with Adapter. For the $\ell$-th transformer layer, the feature update of input token $\mathbf{v}_{i}^{(\ell)}\in\mathbb{R}^{d}$ is computed as:
$\mathbf{v}_{i}^{\prime}=\mathbf{v}_{i}^{(\ell)}+\text{MHSA}(\mathbf{v}_{i}^{(\ell)})$ (28)
$\mathbf{v}_{i}^{a}=\bm{\phi}(\mathbf{v}_{i}^{\prime\top}\mathbf{W}_{\text{down}})\mathbf{W}_{\text{up}}$ (29)
$\mathbf{v}_{i}^{\prime\prime}=\mathbf{v}_{i}^{\prime}+\mathbf{v}_{i}^{a}$ (30)
$\mathbf{v}_{i}^{(\ell+1)}=\mathbf{v}_{i}^{\prime\prime}+\text{FFN}(\mathbf{v}_{i}^{\prime\prime})$ (31)
[…]
where $\bm{\phi}(\cdot)$ is a ReLU activation, $\mathbf{W}_{\text{down}}\in\mathbb{R}^{d\times d_{h}}$ and $\mathbf{W}_{\text{up}}\in\mathbb{R}^{d_{h}\times d}$ are the adapter projection weights, $\mathbf{W}_{\text{proj}}\in\mathbb{R}^{d\times d_{vt}}$ is the projection weight that aligns the video and text feature spaces.
Multi-Head Self-Attention (MHSA). Given token sequence $\mathbf{Z}\in\mathbb{R}^{N\times d}$, multi-head self-attention is computed as:
$\text{MHSA}(\mathbf{Z})=\text{Concat}(\mathbf{h}_{1},\dots,\mathbf{h}_{H})\mathbf{W}^{O},$ (33)
where each head $\mathbf{h}_{h}$ is computed as:
$\mathbf{h}_{h}=\text{Softmax}\left(\frac{\mathbf{Q}_{h}\mathbf{K}_{h}^{\top}}{\sqrt{d_{h}}}\right)\mathbf{V}_{h},$ (34)
with projections:
$\mathbf{Q}_{h}=\mathbf{Z}\mathbf{W}_{h}^{Q},\quad\mathbf{K}_{h}=\mathbf{Z}\mathbf{W}_{h}^{K},\quad\mathbf{V}_{h}=\mathbf{Z}\mathbf{W}_{h}^{V},$ (35)
and projection matrices $\mathbf{W}_{h}^{Q},\mathbf{W}_{h}^{K},\mathbf{W}_{h}^{V}\in\mathbb{R}^{d\times d_{h}}$, and $\mathbf{W}^{O}\in\mathbb{R}^{d\times d}$.
Feedforward Network (FFN). The FFN is a two-layer MLP with GELU activation:
$\text{FFN}(\mathbf{v})=\text{GELU}(\mathbf{v}\mathbf{W}_{1})\mathbf{W}_{2},$ (36)
where $\mathbf{W}_{1}\in\mathbb{R}^{d\times 4d}$ and $\mathbf{W}_{2}\in\mathbb{R}^{4d\times d}$.
Output. After passing through $L$ layers, the final spatial feature of the $i$-th frame is:
$\mathbf{V}_{i}^{s}=\mathcal{F}(V_{i};\mathcal{A}^{b})\in\mathbb{R}^{d_{vt}},\quad i=1,\dots,N_{f}.$ (37)
These frame-level features $\{\mathbf{V}_{i}^{s}\}_{i=1}^{N_{f}}$ are then aggregated by the spatiotemporal encoder to form the global video representation.
B.3 Spatiotemporal Encoder
Let a video clip be uniformly sampled into $N_{f}$ frames. Each frame is processed by a spatial encoder to yield frame features $\mathbf{V}^{s}_{i}$. A learnable [CLS] token $\mathbf{V}^{s}_{\text{cls}}$ is prepended to the sequence, forming the input:
$\tilde{\mathbf{V}}^{s}=[\mathbf{V}^{s}_{\text{cls}};\mathbf{V}^{s}_{1};\cdots;\mathbf{V}^{s}_{N_{f}}].$ (38)
This sequence is passed through a Transformer-based spatiotemporal encoder. Each layer comprises Multi-Head Self-Attention (MHSA), residual connections, and feedforward networks (FFNs). For a single attention head, the attention weights are computed as:
$A_{ij}=\frac{\exp\left(\frac{\langle\mathbf{q}_{i},\mathbf{k}_{j}\rangle}{\sqrt{d_{k}}}\right)}{\sum_{j^{\prime}=1}^{N_{f}+1}\exp\left(\frac{\langle\mathbf{q}_{i},\mathbf{k}_{j^{\prime}}\rangle}{\sqrt{d_{k}}}\right)},$ (39)
where $\mathbf{q}_{i}=\mathbf{W}^{Q}\tilde{\mathbf{V}}^{s}_{i}$, $\mathbf{k}_{j}=\mathbf{W}^{K}\tilde{\mathbf{V}}^{s}_{j}$ and $\mathbf{v}_{j}=\mathbf{W}^{V}\tilde{\mathbf{V}}^{s}_{j}$ are the query, key, and value projections. The output of attention for token $i$ is:
$\mathbf{y}_{i}=\sum_{j=1}^{N_{f}+1}A_{i,j}\mathbf{v}_{j}.$ (40)
We define the final spatiotemporal video representation as the output at the [CLS] token:
$\mathbf{V}^{st}=\mathbf{y}_{\text{cls}}=\sum_{j=1}^{N_{f}+1}A_{\text{cls},j}\mathbf{v}_{j}.$ (41)
This architecture enables spatiotemporal modeling by allowing: (i) long-range temporal interactions across frames via MHSA; (ii) spatial semantics retention from CLIP features; and (iii) adaptive fusion of information through the [CLS] token. Multiple attention heads further enhance expressiveness by learning diverse patterns. Consequently, $\mathbf{V}^{st}$ serves as a content-adaptive spatiotemporal descriptor capturing both motion and appearance.
Appendix C Additional Experiments
C.1 Complexity Analysis: FLOPs and Parameters
We compute the theoretical floating point operations (FLOPs) for each module: the CLIP ViT-B/16 backbone, the inserted adapters, and the spatiotemporal encoder module.
Table 7: FLOPs and parameter counts for each module per 8-frame video.
Module Input Shape Parameters FLOPs (GFLOPs)
CLIP (ViT-B/16) (frozen) $(8,3,224,224)$ 86M 269.81
Adapter $(8,197,768)$ 1.17M (1.36%) 3.73 (1.38%)
Spatiotemporal Encoder $(9,512)$ 9.45M (10.99%) 0.0854 (0.03%)
Table 7 provides a detailed breakdown of computational complexity and parameter count for each component in our framework, evaluated on 8-frame video inputs. The backbone CLIP (ViT-B/16) dominates the overall cost with 86M parameters and 148 GFLOPs. The inserted adapter modules, despite being integrated into every transformer layer, introduce only 1.17M additional parameters (1.36%) and 1.38% more FLOPs, demonstrating their lightweight nature. Furthermore, our spatiotemporal encoder—used to capture dynamic information—adds merely 0.085 GFLOPs (0.03%, per-expert) and 9.45M parameters (10.99%), confirming its computational efficiency. These results validate that our method enhances temporal modeling with minimal overhead, making it well-suited for continual learning in resource-constrained settings.
C.2 Comparison of MoE Routing Strategies (task by task)
Figure 4: MoE Methods Performance per Task on UCF101 and HMDB51
Figure 4 presents the relative accuracy change (%) of different Mixture-of-Experts (MoE) routing strategies on three benchmarks: UCF101 with two task configurations (5$\times$10s and 10$\times$5s), and HMDB51 (5$\times$5s). The horizontal axis represents the incremental task ID, while the vertical axis shows the accuracy change relative to the Avg-MoE baseline.
We observe that TD-MoE consistently outperforms all baselines across datasets and task granularities. Its performance advantage becomes more pronounced as the number of tasks increases, reaching up to 6–8% improvement on later tasks. In contrast, Adapter-MoE and CLIP-MoE exhibit only marginal gains, which tend to saturate early, suggesting limited ability to model task-specific dynamics. These findings confirm that temporal decomposition is effective for guiding expert selection in a task ID-agnostic manner, and helps mitigate forgetting by capturing relevant spatiotemporal cues.
C.3 Effectiveness of Frame-Shared Semantics Distillation (task by task)
Figure 5: Distillation Methods Performance per Task on UCF101 and HMDB51
Figure 5 compares the impact of different distillation strategies, including w/o Distillation, unified Distillation, and our proposed Frame-Shared Semantics Distillation (FSSD), on UCF101 and HMDB51. The vertical axis reports the relative improvement over the non-distillation baseline.
The results demonstrate that FSSD delivers the most consistent and significant improvements, particularly on HMDB51. While unified distillation offers slight improvements, it lacks consistency across tasks. This suggests that uniform constraints fail to address the heterogeneous semantic importance of feature channels. Moreover, the increasing gap between FSSD and other methods over time confirms that adaptive regularization based on frame-shared semantic importance is critical for preserving relevant knowledge across tasks in VCIL.
Appendix D Limitations
One limitation of our approach is that the text classifier is based on a vision-language model pre-trained on natural images and general text corpora. This may introduce a domain gap when applied to specialized video domains (e.g., medical or industrial settings), where the semantic alignment between text and video content may be less reliable. Adapting or fine-tuning the text encoder on domain-specific data is a promising direction for future work.
Appendix E Potential Impact
Our method for video class-incremental learning has potential benefits in areas such as robotics and assistive systems by improving adaptability and efficiency, we acknowledge that any AI-based video understanding system could also be misused for surveillance or privacy-infringing applications. Mitigating such risks requires careful deployment and adherence to ethical standards.