### v2

# StPR: Spatiotemporal Preservation and Routing for Exemplar-Free Video Class-Incremental Learning

Abstract
Video Class-Incremental Learning (VCIL) seeks to develop models that continuously learn new action categories over time without forgetting previously acquired knowledge. Unlike traditional Class-Incremental Learning (CIL), VCIL introduces the added complexity of spatiotemporal structures, making it particularly challenging to mitigate catastrophic forgetting while effectively capturing both frame-shared semantics and temporal dynamics. Existing approaches either rely on exemplar rehearsal, raising concerns over memory and privacy, or adapt static image-based methods that neglect temporal modeling. To address these limitations, we propose Spatiotemporal Preservation and Routing (StPR) mechanism, a unified and exemplar-free VCIL framework that explicitly disentangles and preserves spatiotemporal information. We begin by introducing Frame-Shared Semantics Distillation (FSSD), which identifies semantically stable and meaningful channels by jointly considering channel-wise sensitivity and classification contribution. By selectively regularizing these important semantic channels, FSSD preserves prior knowledge while allowing for adaptation. Building on this preserved semantic space, we further design a Temporal Decomposition-based Mixture-of-Experts (TD-MoE), which dynamically routes task-specific experts according to temporal dynamics, thereby enabling inference without task IDs or stored exemplars. Through the synergy of FSSD and TD-MoE, StPR progressively leverages spatial semantics and temporal dynamics, culminating in a unified, exemplar-free VCIL framework. Extensive experiments on UCF101, HMDB51, SSv2 and Kinetics400 show that our method outperforms existing baselines while offering improved interpretability and efficiency in VCIL.
1 Introduction
Class-Incremental Learning (CIL) li2017learning ; belouadah2021comprehensive ; de2021continual ; masana2022class ; zhang2024center develops models that learn from a sequence of tasks without forgetting previous knowledge, recognizing an ever-growing set of classes without past task data or identifiers. A key challenge is catastrophic forgetting CatastrophicInterference1989McCloskey ; ConnectionistModels1990Ratcliff , where new knowledge overwrites old. While well studied for images, extending CIL to videos: Video Class-Incremental Learning (VCIL) park2021class ; villa2022vclimb , remains underexplored. VCIL differs from CIL by requiring continual learning of new categories while modeling frame-shared semantics and temporal dependencies, unlike CIL’s focus on static images. This spatiotemporal complexity is critical for understanding actions, motion, and scene dynamics in real-world applications like surveillance, driver monitoring, and robotics. Further, memory and privacy constraints often prohibit storing past data, demanding continual learning without rehearsal.
Figure 1: (a) Prior methods rely on exemplar rehearsal or naively stack video and CIL modules. (b) Our StPR framework explicitly decouples and reuses spatiotemporal semantics to mitigate forgetting.
The central challenge of VCIL lies in mitigating catastrophic forgetting while effectively leveraging frame-shared semantics and temporal dynamics to incrementally learn new categories. Existing methods can be broadly categorized into two types, as illustrated in Figure 1(a): 1) Exemplar-based methods rebuffi2017icarl ; hou2019learning ; douillard2020podnet ; park2021class ; pei2022learning ; villa2022vclimb ; alssum2023just ; liang2024hypercorrelation ; chen2025csta store a portion of previous data (video clips, frames, or features) and apply rehearsal to reduce forgetting. However, storing exemplars incurs memory and privacy costs and typically emphasizes frame-level learning without explicitly modeling temporal dynamics. 2) CIL-based methods li2017learning ; dhar2019learning ; cheng2024stsp adapt algorithms developed for static images, using techniques like regularization or subspace projection. While avoiding exemplar storage, they often overlook spatiotemporal properties by flattening or underutilizing temporal features. In contrast, our method StPR (Figure 1(b)) explicitly decouples video features into Frame shared semantics and temporal dynamics, and reuses these decomposed components to enhance the model’s ability to adapt continually, thereby reducing forgetting without storing extensive exemplars.
Specifically, we propose a unified, exemplar-free VCIL framework that fully exploits the spatiotemporal nature of videos. Our method integrates both spatial semantic consistency and temporal variation to mitigate forgetting and improve adaptation across tasks. Separately, we introduce: 1) Frame-Shared Semantics Distillation (FSSD). To preserve frame-shared semantics and alleviate forgetting, we quantify the semantic importance of each channel using a combination of semantic sensitivity and classification contribution. This ensures that semantically meaningful and stable channels are preserved, achieving a better trade-off between stability and plasticity. 2) Temporal-Decomposition-based Mixture-of-Experts (TD-MoE). To exploit temporal dynamics for continual adaptation, we decouple task-specific temporal cues for each expert. At inference, expert routing depends solely on the temporal dynamics of the input, without requiring task identities or stored exemplars. This enables dynamic assignment of weights to experts according to the temporal dynamics of the input, facilitating incremental learning of new categories.
Our framework uniquely bridges the gap between video-specific spatiotemporal representation and class-incremental adaptation. By disentangling and leveraging both spatial semantic channel consistency and temporal dynamics, it offers an effective and explainable solution for continual video understanding. Our main contributions are: 1) We propose a Frame-Shared Semantics Distillation method (FSSD) that preserves frame-shared, semantically aligned spatial channels through semantic importance-aware regularization, optimizing the stability-plasticity trade-off in continual learningl; 2) We design a Temporal Decomposition based Mixture-of-Experts strategy (TD-MoE) that decomposes spatiotemporal features and uses temporal dynamics for expert combination, enabling task-id-free and dynamic adaptation; 3) We present a unified, exemplar-free VCIL framework that achieves state-of-the-art results on UCF101, HMDB51, SSv2 and Kinetics400, demonstrating the effectiveness of integrating spatial semantics and temporal dynamics in VCIL.
2 Related Work
2.1 Class-Incremental Learning
Class-Incremental Learning (CIL) aims to enable models to continually learn new classes without forgetting previously learned ones. Existing approaches typically fall into three categories: (1) regularization-based methods kirkpatrick2017overcoming ; zenke2017continual ; xiang2022coarse ; zhou2023hierarchical , which constrain parameter updates to preserve prior knowledge, often via knowledge distillation li2017learning ; hou2019learning ; (2) exemplar-based methods bang2021rainbow ; chaudhry2018riemannian ; rebuffi2017icarl , which store or generate past data to reduce forgetting; and (3) structure-based methods serra2018overcoming ; mallya2018packnet ; mallya2018piggyback ; liang2024inflora ; yu2024boosting , which expand model capacity or isolate task-specific components. Recently, CIL combined with pre-trained vision transformers (ViTs) ermis2022memory ; CODAPromptCOntinual2023Smith ; wang2022dualprompt ; wang2022learning has achieved impressive results by leveraging transferable representations and modularity. Some methods fully fine-tune pre-trained backbones boschini2022transfer ; zhang2023slca , but this is computationally expensive. To address efficiency, parameter-efficient fine-tuning (PEFT) methods have been introduced. Prompt pool-based approaches wang2022learning ; CODAPromptCOntinual2023Smith ; wang2024hierarchical ; zhang2023slca maintain task-specific prompts, while adapter-based methods zhou2023revisiting ; tan2024semantically ; gao2024beyond ; liang2024inflora ; ExpandableSubspace2024Zhou adapt ViTs to new classes with minimal updates. While effective, most CIL strategies were originally developed for static image domains and do not generalize well to video-based scenarios, where temporal dynamics play a critical role.
2.2 Video Class-Incremental Learning
Action recognition has been widely explored with 2D CNNs using temporal aggregation lin2019tsm ; wang2016temporal and 3D CNNs for joint spatiotemporal modeling carreira2017quo ; tran2015learning . More recent work focuses on improving temporal sensitivity and efficiency feichtenhofer2020x3d ; fan2020rubiksnet . However, these models are trained in static setups and do not address continual adaptation or forgetting. To address these challenges, Video Class-Incremental Learning (VCIL) extends conventional Class-Incremental Learning (CIL) to spatiotemporal data, introducing additional challenges such as managing temporal variations across tasks. Several recent methods, including TCD park2021class , FrameMaker pei2022learning , and HCE liang2024hypercorrelation , address this setting by storing videos or compressed exemplars. However, these strategies raise concerns related to memory efficiency and data privacy. While SMILE alssum2023just effectively extracts image features from individual frames, it does not explicitly capture temporal information, which may limit its ability to leverage the distinctive decision cues present in video data. Exemplar-free methods such as STSP cheng2024stsp mitigate forgetting via orthogonal subspace projections, but they mainly adapt image-domain strategies to video tasks. In contrast, our approach decouples and models the spatiotemporal structure of videos, proposing a unified VCIL framework that preserves spatial consistency via Frame-Shared Semantics Distillation (FSSD) for knowledge retention without exemplars, while leveraging temporal dynamics for expert routing through Temporal Decomposition-based Mixture-of-Experts (TD-MoE).
Figure 2: Overview of our proposed framework. (I) During training, the (a) Frame-Shared Semantics Distillation (FSSD) module retains past knowledge via frame-shared semantics importance-aware regularization. (II) At inference, the (b) Temporal Decomposition-based Mixture-of-Experts (TD-MoE) dynamically routes input videos to expert branches by leveraging the disentangled temporal component of the spatiotemporal representation, enabling adaptive predictions.
3 Method
Problem Definition: In the Video Class-Incremental Learning (VCIL) setting, a model is trained across $B$ stages with sequentially arriving datasets $\{\mathcal{D}^{1},\dots,\mathcal{D}^{B}\}$. Each dataset $\mathcal{D}^{b}=\{(V_{j}^{b},y_{j}^{b})\}_{j=1}^{|\mathcal{D}^{b}|}$ corresponds to the $b$-th task, where $V_{j}^{b}$ is the $j$-th video and $y_{j}^{b}$ is its class label. Here, videos primarily represent human action recognition scenarios, where the spatiotemporal dynamics capture motion patterns, subject interactions, and scene context. $|\mathcal{D}^{b}|$ represents the number of samples in the $b$-th task. Let $\mathcal{Y}^{b}$ be the label space of the $b$-th dataset. For all $b\neq b^{\prime}$, the label spaces are disjoint: $\mathcal{Y}^{b}\cap\mathcal{Y}^{b^{\prime}}=\varnothing$. The objective of VCIL is to incrementally train a model over $B$ tasks while maintaining high performance across all accumulated classes $\{\mathcal{Y}^{1},\mathcal{Y}^{2},\dots,\mathcal{Y}^{B}\}$.
Overall framework. We propose a unified, exemplar-free framework for Video Class-Incremental Learning (VCIL) built upon the CLIP model radford2021learning . Our goal is to mitigate catastrophic forgetting while effectively leveraging frame-shared semantics and temporal dynamics to incrementally learn new categories. The frozen visual encoder $\mathcal{F}(\cdot)$ extracts spatial features, while adapters $\mathcal{A}^{b}$ are updated for each task $b$. A spatiotemporal encoder $\mathcal{G}(\cdot)$ models temporal dynamics. Our framework introduces two key components: 1) Frame-Shared Semantics Distillation (FSSD) identifies semantically stable channels across frames by combining Semantic Sensitivity and Classification Score, applying selective regularization to preserve critical spatial semantics while maintaining plasticity. 2) Temporal Decomposition based Mixture-of-Experts (TD-MoE). To exploit temporal dynamics for continual adaptation, we decouple shared static components and temporal dynamics. During inference, temporal dynamics are used to assign dynamic weights to expert temporal encoders, enabling task-id-free adaptation without requiring task identifiers or stored exemplars.
3.1 Spatial and Spatiotemporal Encoder
Spatial Encoder. The shared adapter module chen2022adaptformer $\mathcal{A}^{b}=\{\mathcal{A}_{l}^{b}\}_{l=1}^{N}$ is utilized with a frozen CLIP-ViT model with $N$ layers of transformer module, serving as the spatial extractor. An adapter is an encoder-decoder architecture embedded into the residual of each transformer layer, facilitating transfer learning and enhancing downstream task performance. Typically, it consists of a down-sampling MLP $\mathbf{W}_{down}\in\mathbb{R}^{d\times d_{h}}$, a ReLU activation $\bm{\phi}(\cdot)$, and an up-sampling MLP $\mathbf{W}_{up}\in\mathbb{R}^{d_{h}\times d}$, where $d$ is the input and output dimension, and $d_{h}$ is the hidden dimension. For adapter, if input is $\mathbf{v}_{i}^{\prime}\in\mathbb{R}^{d}$, which is the $i$-th frame-level feature after the Multi-Head Self-Attention and residual connection in CLIP ViT, the output of adapter is:
$\mathbf{v}^{a}_{i}=\bm{\phi}(\mathbf{v}_{i}^{\prime}\mathbf{W}_{down})\mathbf{W}_{up}.$ (1)
The spatial feature of $i$-th frame $\mathbf{V}_{i}^{s}=\mathcal{F}(V_{i};\mathcal{A}^{b}),\mathbf{V}_{i}^{s}\in\mathbb{R}^{d_{vt}}$, where $V_{i}$ is the $i$-th frame of the video, $d_{vt}$ is the dimension of the aligned video-text features.
Spatiotemporal Encoder. To obtain the spatiotemporal representation, we feed both the frame-level spatial features and a learnable [CLS] token into a multi-head self-attention based spatiotemporal encoder $\mathcal{G}(\cdot)$. Specifically, the input to $\mathcal{G}$ consists of the frame features $\mathbf{V}^{s}_{1},\dots,\mathbf{V}^{s}_{N_{f}}$ and a [CLS] token $\mathbf{V}^{s}_{cls}\in\mathbb{R}^{d_{vt}}$, where $N_{f}$ represents the number of sampled frames. The temporal encoder outputs the spatiotemporal feature $\mathbf{V}^{st}$, corresponding to the transformed [CLS] token:
$\mathbf{V}^{st}=\mathcal{G}([\mathbf{V}^{s}_{cls};\mathbf{V}^{s}_{1};\dots;\mathbf{V}^{s}_{N_{f}}])[0]\in\mathbb{R}^{d_{vt}},$ (2)
where $[0]$ selects the output associated with the [CLS] token after attention-based aggregation.
3.2 Frame-Shared Semantics Distillation
In VCIL, shared adapter modules inevitably drift in feature space when adapting to new tasks, leading to forgetting. Directly applying classic uniform-weighted distillation from CIL to video tasks ignores the differences in semantic importance and temporal stability across video features. This leads to a suboptimal balance between stability and plasticity. To address this, we propose Frame-Shared Semantics Distillation (FSSD), which identifies stable cross-frame channels capturing core semantics and regularizes them to preserve prior knowledge while allowing adaptation
Frame-Shared Semantics Distillation (FSSD). To mitigate semantic drift across tasks, we introduce a distillation loss weighted by the frame-shared semantics importance:
$\mathcal{L}_{FSSD}=\frac{1}{|\mathcal{D}^{b}|\cdot d_{vt}}\sum_{c}^{|\mathcal{C}_{b}|}\sum_{i}^{N_{c}}\sum_{j}^{d_{vt}}I_{b-1,c,j}\cdot\|\bar{V}_{b-1,c,i,j}^{s}-\bar{V}_{b,c,i,j}^{s}\|_{2}^{2},$ (3)
where $|\mathcal{C}_{b}|$ is the number of classes in the $b$-th task, and $N_{c}$ is the number of samples per class. $I_{b-1,c,j}$ denotes the frame-shared importance of the $j$-th channel for class $c$ from task (b1)(b{-}1). $\bar{V}_{b-1,c,i,j}^{s}$ and $\bar{V}_{b,c,i,j}^{s}$ are the $j$-th channel outputs of the $i$-th sample in class $c$, extracted from the spatial encoders of task (b1)(b{-}1) and $b$, respectively, calculated on current data, with the previous model frozen.
Frame-Shared Semantics. To quantify the importance of frame-shared semantics, we assess each channel based on two criteria: 1). Semantic Sensitivity. It measures the responsiveness to activation changes, thereby reflecting its reliability in representing consistent semantic information. and 2) Classification Score. It reflects the channel’s contribution to the final classification.
For semantic sensitivity, we employ Fisher Information to estimate how sensitively a channel’s activation influences the output. As spatial features $\bar{\mathbf{V}}^{s}=\frac{1}{N_{f}}\sum_{i=1}^{N_{f}}\mathbf{V}^{s}_{i}$ aggregate frame-wise variations, the Central Limit Theorem suggests each channel’s distribution approximates a Gaussian (Belong to the same category). Thus, we assume the $j$-th channel activation for class $c$ follows (For simplicity, we omit the subscripts for task and sample.):
$\bar{V}^{s}_{c,j}\sim\mathcal{N}(\mu_{c,j},\sigma_{c,j}^{2}),$ (4)
where $\mu_{c,j}$ and $\sigma_{c,j}^{2}$ denote the mean and variance across frames. The Fisher Information $\mathcal{I}(\mu_{c,j})$ (Detailed derivations are provided in the appendix B) with respect to $\mu_{c,j}$ is computed as:
$\mathcal{I}(\mu_{c,j})=\mathbb{E}\left[\left(\frac{\bar{V}^{s}_{c,j}-\mu_{c,j}}{\sigma_{c,j}^{2}}\right)^{2}\right]=\frac{1}{\sigma_{c,j}^{4}}\mathbb{E}\left[(\bar{V}^{s}_{c,j}-\mu_{c,j})^{2}\right]=\frac{1}{\sigma_{c,j}^{4}}\cdot\sigma_{c,j}^{2}=\frac{1}{\sigma_{c,j}^{2}}.$ (5)
For classification score, we compute the cosine similarity between the spatial video feature $\bar{\mathbf{V}}^{s}_{c}\in\mathbb{R}^{d_{vt}}$ and its corresponding text feature $\mathbf{T}_{c}\in\mathbb{R}^{d_{vt}}$. Specifically, for the $j$-th channel, the classification score is defined as $\gamma_{c,j}=\frac{\bar{V}^{s}_{c,j}\cdot T_{c,j}}{\|\bar{\mathbf{V}}^{s}_{c}\|\cdot\|\mathbf{T}_{c}\|}$, where $T_{c,j}$ denotes the $j$-th dimension feature of $\mathbf{T}_{c}$. We then take the expectation of $\gamma_{c,j}$ across frames to obtain a stable channel-level contribution estimate:
$\mathbb{E}[\gamma_{c,j}]=\mathbb{E}\left[\frac{\bar{V}^{s}_{c,j}\cdot T_{c,j}}{\|\bar{\mathbf{V}}^{s}_{c}\|\cdot\|\mathbf{T}_{c}\|}\right]\propto\mathbb{E}\left[\frac{\bar{V}^{s}_{c,j}\cdot T_{c,j}}{\|\bar{\mathbf{V}}^{s}_{c}\|}\right]\approx\frac{T_{c,j}\cdot\mu_{c,j}}{\lambda},$ (6)
where $\|\bar{\mathbf{V}}^{s}_{c}\|\approx\lambda$ is treated as a constant after normalization. Combining semantic sensitivity and classification score, the semantic importance for the $j$-th channel of the $c$-th class is defined as:
$I_{c,j}=\frac{T_{c,j}\cdot\mu_{c,j}}{\sigma_{c,j}^{2}}.$ (7)
FSSD accumulates frame-shared semantic importance as distillation weights, retaining key channels for old tasks while allowing less important ones to adapt, thus balancing stability and plasticity.
3.3 Temporal Decomposition based Mixture-of-Experts
Given the high forgetting tendency of deep transformers in VCIL, we allocate a dedicated spatiotemporal encoder for each task. As task IDs are unavailable during inference, we allocate a spatiotemporal encoder per task and design a routing mechanism that dynamically weights experts based on temporal patterns, ensuring relevant experts contribute more to the final representation.
Task-Specific Expert. For each task, we train a dedicated expert based on the spatiotemporal encoder. The spatiotemporal features $\mathbf{V}^{st}\in\mathbb{R}^{d_{vt}}$ captured by each expert are computed as in Eq. 2.
Temporal Decomposition-based Router.
To design this routing mechanism based on temporal dynamics, we consider two aspects: 1) Temporal residuals. These reflect the subtle temporal differences within redundant frames. 2) Inter-frame information. This captures abstract temporal concepts between frames, based on the knowledge learned by each expert.
Figure 3: For each video, we uniformly sample 8 frames to compute the $p$-value defining $p>0.05$ as strictly stationary and $p>0.02$ as weakly stationary in the short term.
For temporal residuals, we observe that redundant frames, where backgrounds and subjects remain consistent, cause minimal variation between adjacent frames kim2024fusion ; liu2021no . This leads to short-term temporal stationarity, which we further validate on the UCF101 and HMDB51 datasets Fig. 3. Thus, each frame feature is decomposed as $\mathbf{V}_{i}^{s}=\bar{\mathbf{v}}+\bm{\epsilon}_{i}$, with $\bar{\mathbf{v}}$ as shared static components and $\bm{\epsilon}_{i}$ as temporal residuals. The spatial representation is then the mean across frames:
$\bar{\mathbf{V}}^{s}=\bar{\mathbf{v}}+\bar{\bm{\epsilon}},\quad\bar{\bm{\epsilon}}=\frac{1}{N_{f}}\sum_{i=1}^{N_{f}}\bm{\epsilon}_{i}.$ (8)
For the inter-frame information, since the spatiotemporal feature $\mathbf{V}^{st}$ is computed by the attention module, it can be approximated as $\mathbf{V}^{st}\approx\sum_{i=1}^{N_{f}}a_{i}\cdot\mathbf{V}^{s}_{i}$, where $a_{i}$ is attention score. After normalization, we can obtain $\sum_{i=1}^{N_{f}}a_{i}=1$. Substituting Eq. 8, we obtain:
$\mathbf{V}^{st}=\bar{\mathbf{v}}+\sum_{i=1}^{N_{f}}a_{i}\cdot\bm{\epsilon}_{i}.$ (9)
Since $\bar{\mathbf{v}}$ is difficult to estimate, and to decouple the temporal residual $\bm{\epsilon}_{i}$ and inter-frame information $a_{i}$, we naturally address this by using the difference between $\mathbf{V}^{st}$ and $\bar{\mathbf{V}}^{s}$, effectively isolating the temporal dynamics, which can be represented as:
$\mathbf{V}^{tem}=\sum_{i=1}^{N_{f}}\left(a_{i}-\tfrac{1}{N_{f}}\right)\cdot\bm{\epsilon}_{i}.$ (10)
This formulation reveals that $\mathbf{V}^{tem}$ quantifies the deviation between the model’s attention-weighted temporal dynamics and the uniform temporal mean, effectively disentangling temporal variations from static semantics. This enables routing to exploit temporal cues while avoiding background interference, thereby mitigating forgetting and enhancing continual learning.
Inference. During inference, we first compute the decoupled temporal representation $\mathbf{V}^{tem}\in\mathbb{R}^{d_{vt}}$ for each input video. For all categories in the current task, we calculate the mean temporal representation and store it in the anchor pool as $\bar{\mathbf{V}}_{c}^{tem}\in\mathbb{R}^{d_{vt}}$, where $c$ represents the $c$-th class. For each expert $k$, we compute a similarity-based score as the router:
$r_{k}=\max_{c\in\mathcal{C}_{k}}\cos\left(\mathbf{V}_{k}^{tem},\bar{\mathbf{V}}_{c}^{tem}\right),$ (11)
where $\mathcal{C}_{k}$ represents the set of classes assigned to expert $k$. Then, we combine the adapter-tuned spatial features $\bar{\mathbf{V}}^{s}$ with the expert outputs weighted by $r_{k}$ as the final video representation:
$\mathbf{V}=\bar{\mathbf{V}}^{s}+\sum_{k}r_{k}\cdot\mathbf{V}_{k}^{st}.$ (12)
The final video representation is matched with text embedding via cosine similarity for classification.
3.4 Loss Function and Optimization
Our loss function includes: 1) contrastive loss between video features and text descriptions for classification; 2) contrast loss between video features under adapter fine-tuning and text features for spatial optimization; and 3) FSSD loss to mitigate forgetting in shared adapter modules.
Contrastive Loss Formulation. We use symmetric contrastive loss for video-to-text and text-to-video alignment. Given a batch of $N$ samples, let $\mathbf{V}_{i}$ and $\mathbf{T}_{j}$ denote the video and text features, respectively. The similarity between video $i$ and text $j$ is computed as the cosine similarity $S_{i,j}=\cos(\mathbf{V}_{i},\mathbf{T}_{j})$, forming a similarity matrix $S\in\mathbb{R}^{N\times N}$. Let $\mathbf{M}\in\{0,1\}^{N\times N}$ be the label mask, where $M_{i,j}=1$ if $y_{i}=y_{j}$ and $M_{i,j}=0$ otherwise. Then, the Video-to-text contrastive loss can be caculated by:
$\mathcal{L}_{\text{v2t}}=-\frac{1}{N}\sum_{i=1}^{N}\log\left(\frac{\sum_{j=1}^{N}M_{i,j}\cdot\exp(S_{i,j})}{\sum_{j=1}^{N}\exp(S_{i,j})+\varepsilon}\right),$ (13)
where $\varepsilon$ is a small constant added to avoid division by zero. The Text-to-Video loss $\mathcal{L}_{\text{t2v}}$ is similarly defined by swapping video and text in the equation. Symmetric total contrastive loss is :
$\mathcal{L}_{\text{Cont}}=\frac{1}{2}(\mathcal{L}_{\text{v2t}}+\mathcal{L}_{\text{t2v}}).$ (14)
For $\mathcal{L}_{\text{Cont}}^{\text{St}}$, the embeddings are the spatiotemporal video feature $\mathbf{V}^{st}$ and corresponding text features. For $\mathcal{L}_{\text{Cont}}^{\text{S}}$, we use the CLIP adapter feature $\bar{\mathbf{V}}^{s}$ and corresponding text features.
Total Loss. The overall training loss is:
$\mathcal{L}=\mathcal{L}_{\text{Cont}}^{\text{St}}+\mathcal{L}_{\text{Cont}}^{\text{S}}+w\cdot\mathcal{L}_{\text{FSSD}},$ (15)
where $w$ is a hyperparameter. This design aligns both spatial and spatiotemporal semantics with text supervision, while the FSSD term preserves critical frame-shared semantics to mitigate forgetting.
4 Experiments
Table 1: Average Accuracy ($\overline{\mathrm{Acc}}$) of the UCF101 and HMDB51 under the TCD benchmark.
Method Exemplar Venue UCF101 HMDB51
$10\times 5$s $5\times 10$s $2\times 25$s $5\times 5$s $1\times 25$s
iCaRL rebuffi2017icarl CVPR’17 $65.34$ $64.51$ $58.73$ $40.09$ $33.77$
LwFMC li2017learning TPAMI’18 $42.14$ $25.59$ $11.68$ $26.82$ $16.49$
LwM dhar2019learning CVPR’19 $43.39$ $26.07$ $12.08$ $26.97$ $16.50$
UCIR hou2019learning CVPR’19 $74.09$ $70.50$ $64.00$ $46.53$ $37.15$
PODNet douillard2020podnet ECCV’20 $74.37$ $73.75$ $71.87$ $48.78$ $46.62$
TCD park2021class ICCV’21 $77.16$ $75.35$ $74.01$ $50.36$ $46.66$
FrameMaker pei2022learning NeurIPS’22 $78.64$ $78.14$ $77.49$ $51.12$ $47.37$
L2P wang2022learning CVPR’22 $81.24$ $80.09$ $78.58$ $49.98$ $45.87$
S-iPrompts wang2022s NeurIPS’22 $80.60$ $80.27$ $80.43$ $53.11$ $53.89$
ST-Prompt†pei2023space CVPR’23 $84.75$ $85.54$ $85.67$ $60.14$ $60.54$
STSP cheng2024stsp ECCV’24 $81.15$ $82.84$ $79.25$ $56.99$ $49.19$
HCE liang2024hypercorrelation AAAI’24 $80.01$ $78.81$ $77.62$ $52.01$ $48.94$
StPR (Ours) $\mathbf{-}$ $\mathbf{94.67}$ $\mathbf{92.13}$ $\mathbf{88.52}$ $\mathbf{68.12}$ $\mathbf{67.01}$
Table 2: Average Accuracy ($\overline{\mathrm{Acc}}$) of the SSv2 under the TCD benchmark, with best results in bold.
Method Exemplar Venue $\textbf{10}\times\textbf{9}$s $\textbf{5}\times\textbf{18}$s
iCaRL rebuffi2017icarl CVPR’17 $20.41$ $16.62$
UCIR hou2019learning CVPR’19 $24.32$ $19.31$
PODNet douillard2020podnet ECCV’20 $27.63$ $20.14$
TCD park2021class ICCV’21 $29.32$ $24.69$
FrameMaker pei2022learning NeurIPS’22 $31.41$ $26.57$
L2P wang2022learning CVPR’22 $26.02$ $21.33$
S-iPrompts wang2022s NeurIPS’22 $33.69$ $30.84$
ST-Prompt†pei2023space CVPR’23 $39.98$ $35.44$
HCE liang2024hypercorrelation AAAI’24 $36.88$ $32.82$
StPR (Ours) $\mathbf{-}$ $\mathbf{40.79}$ $\mathbf{37.30}$
Table 3: Results of the Kinetics-400 under the vCLIMB benchmark at 10 and 20 task settings.
Method Exemplars Venue Kinetics400-10s Kinetics400-20s
[…]
4.1 Experimental Details
Dataset. We evaluate our method on four benchmark datasets: UCF101 soomro2012ucf101 , HMDB51kuehne2011hmdb , Something-Something V2 (SSv2) goyal2017something and Kinetics400 carreira2017quo . All experiments are conducted in an exemplar-free setting. For fair comparison, we use the TCD benchmark park2021class on UCF101, HMDB51, and SSv2, pretraining the model on 51, 26, and 84 base classes, respectively, with the remaining classes split into tasks. For Kinetics-400, we follow the vCLIMB benchmark villa2022vclimb with 10- or 20-task splits, each containing the same number of classes.
Evaluation Metrics. We adopt three widely-used metrics to evaluate performance in VCIL: 1). Final Accuracy (Acc) villa2022vclimb , which measures the overall classification accuracy on all learned classes after the final task is completed; 2). Average Accuracy ($\overline{\mathrm{Acc}}$) park2021class , which measures the mean classification accuracy over all incremental stages after the final task is completed; 3). Backward Forgetting (BWF) villa2022vclimb , which quantifies the average drop in performance on previously learned tasks, reflecting how well the model retains past knowledge.
Implementation Details. All experiments are conducted on a single NVIDIA RTX 3090 GPU. We adopt the CLIP ViT-B/16 model radford2021learning as the backbone, with all its parameters frozen during training. The spatial and spatiotemporal encoders are the only trainable components in our framework. For optimization, we employ Stochastic Gradient Descent (SGD) with an initial learning rate of 0.01 and a batch size of 40. Each task is trained for 60 epochs in the first incremental session and 30 epochs in each subsequent session. The weighting hyperparameter $w$ in Eq. 15 is set to $1\times 10^{4}$. The multi-head self-attention module within the spatiotemporal encoder consists of three transformer layers, each employing two attention heads. Video clips are sampled using the TSN strategy wang2018temporal , selecting 8 frames per video uniformly across the temporal dimension.
4.2 Main Results
Table. 1, 2 and 3 report results on UCF101, HMDB51, SSv2 and Kinetics400, covering different action complexities and temporal dynamics. Based on their strategies to mitigate forgetting, existing methods are categorized into two groups: 1) Exemplar-based methods (iCaRL, UCIR, PODNet, TCD, FrameMaker, HCE, vCLIMB, SMILE, CSTA). They store video clips, frames, or compressed features and apply rehearsal to reduce forgetting. However, these methods face scalability and privacy challenges due to their reliance on stored exemplars. 2) CIL-based methods (LwFMC, LwM, L2P, S-iPrompts, ST-Prompt†, STSP). This group adapts techniques from image-based class-incremental learning, such as unified distillation and subspace projection, without storing exemplars. While avoiding exemplar storage, their performance tends to be lower, especially as task difficulty increases and lacking explainable spatiotemporal disentanglement. In contrast, Our method (StPR) without storing exemplars, surpasses all baselines across datasets and settings. On the TCD benchmark, our method outperforms the state-of-the-art approach (ST-Prompt†) as well as all exemplar-based methods on UCF101, HMDB51, and SSv2. On the vCLIMB benchmark, exemplar-based methods can alleviate forgetting by replaying stored samples, which makes forgetting lower. Nevertheless, our method achieves higher final accuracy, surpassing the current state-of-the-art (CSTA) and all exemplar-based counterparts.
Table 4: Ablation Study on UCF101 and HMDB51, with best results in bold.
Idx $\mathcal{A}^{b}$ FSSD TD-MoE UCF101($5\times 10$s) HMDB51($5\times 5$s) UCF101($10\times 5$s) HMDB51($25\times 1$s)
Acc $\uparrow$ $\overline{\mathrm{Acc}}\uparrow$ BWF $\downarrow$ Acc $\uparrow$ $\overline{\mathrm{Acc}}\uparrow$ BWF $\downarrow$ Acc $\uparrow$ $\overline{\mathrm{Acc}}\uparrow$ BWF $\downarrow$ Acc $\uparrow$ $\overline{\mathrm{Acc}}\uparrow$ BWF $\downarrow$
1 70.65 74.67 7.01 43.30 43.62 $\mathbf{6.18}$ 70.14 72.72 $5.33$ 43.71 47.48 8.74
2 77.55 81.84 5.76 53.23 55.67 7.73 77.63 82.06 $\mathbf{4.86}$ 54.63 60.83 9.01
3 79.33 89.36 12.38 56.12 61.14 11.75 85.94 93.47 7.40 62.54 68.88 10.20
4 83.07 91.28 10.52 57.47 63.37 21.30 88.03 94.14 8.39 64.71 73.02 21.72
5 $\mathbf{85.79}$ $\mathbf{92.13}$ $\mathbf{5.63}$ $\mathbf{63.04}$ $\mathbf{68.12}$ 11.04 $\mathbf{88.85}$ $\mathbf{94.67}$ $6.31$ $\mathbf{69.61}$ $\mathbf{75.07}$ $\mathbf{7.02}$
4.3 Ablation Study
We perform ablation studies to evaluate the contribution of each component: the adapter tuning ($\mathcal{A}^{b}$), Frame-Shared Semantics Distillation (FSSD), and Temporal Decomposition-based Mixture-of-Experts (TD-MoE). Results are summarized in Table 4. The baseline (pretrained CLIP) model exhibits limited performance on downstream tasks, as it lacks adaptation to new task-specific categories. Introducing FSSD alone moderately improves performance by preserving spatial semantics and reducing semantic drift, while TD-MoE independently enhances adaptation by leveraging temporal dynamics. However, using either module alone yields suboptimal performance. Combining adapter tuning with TD-MoE provides further improvements but still lacks sufficient stability in preserving spatial semantics. The full model (StPR), integrating both FSSD and TD-MoE, achieves the most stable performance across tasks, demonstrating the complementary strengths of spatial semantic preservation and temporal dynamic modeling.
4.4 Further analysis
Analysis of temporal-decomposition routing strategies. Table 5 compares our TD-MoE with several alternative Mixture-of-Experts (MoE) jacobs1991adaptive routing strategies. Simple averaging (Avg-MoE) and static weight assignments—including CLIP-MoE, which uses frozen CLIP visual features for routing, and Adapter-MoE, which uses adapter-tuned CLIP features—provide moderate improvements but fail to dynamically leverage task-specific temporal cues, often resulting in higher forgetting. In contrast, TD-MoE enables adaptive expert weighting based on temporal dynamics, consistently improving both accuracy and stability across tasks. This highlights the importance of modeling temporal variability explicitly, rather than relying on static or feature-agnostic routing.
Table 5: MoE Method Results on UCF101 and HMDB51, with best results in bold.
Method UCF101($5\times 10$s) HMDB51($5\times 5$s) UCF101($10\times 5$s) HMDB51($25\times 1$s)
[…]
Table 6: Distillation Method Results on UCF101 and HMDB51, with best results in bold.
Method UCF101($5\times 10$s) HMDB51($5\times 5$s) UCF101($10\times 5$s) HMDB51($25\times 1$s)
[…]
Effectiveness of FSSD over Uniform Distillation. Table 6 compares our FSSD method with the no-distillation baseline (w/o Distillation) and standard uniform distillation (Distillation) across four VCIL settings. While uniform distillation improves accuracy and reduces backward forgetting (BWF) over the naive baseline, FSSD consistently outperforms both, achieving the highest accuracy and lowest BWF in all settings. These results highlight the benefit of selectively preserving frame-shared semantics, validating the importance-aware design of FSSD for continual video learning. For more experiments (such as hyperparameter analysis and visualization), see the appendix C
5 Conclusion
In this work, we propose StPR, a unified and exemplar-free framework for Video Class-Incremental Learning (VCIL) to tackle the spatiotemporal challenges in continual video learning. By disentangling spatial semantics and temporal dynamics, StPR effectively balances stability and plasticity without relying on stored exemplars. Our method combines Frame-Shared Semantics Distillation (FSSD), which selectively preserves meaningful and stable semantic channels, protecting model’s plasticity. Temporal-Decomposition-based Mixture-of-Experts (TD-MoE), adaptively routes inputs based on temporal cues, reducing forgetting in deep networks. Extensive experiments on UCF101, HMDB51, and SSV2 validate the effectiveness and efficiency of our approach, establishing new state-of-the-art results for continual video recognition. In future work, we plan to explore more realistic application scenarios, such as open-world settings, and investigate the deployment of our method on resource-constrained edge devices.
Appendix
Appendix A Algorithm
Input: Current task data $\mathcal{D}_{b}$; frozen model from task b1b{-}1; text features $\{\mathbf{T}_{c}\}$
Output: FSSD loss $\mathcal{L}_{\text{FSSD}}$
3pt
for each class $c\in\mathcal{Y}_{b}$ do
   Compute mean spatial features:
 $\bar{\mathbf{V}}^{s}_{b,c}=\frac{1}{N_{f}}\sum_{i=1}^{N_{f}}\mathbf{V}^{s}_{b,c,i}$ ;
 // Current model
 $\bar{\mathbf{V}}^{s}_{b-1,c}=\frac{1}{N_{f}}\sum_{i=1}^{N_{f}}\mathbf{V}^{s}_{b-1,c,i}$ ;
 // Previous model
 for each channel $j=1$ to $d$ do
      Estimate $\mu_{c,j},\sigma^{2}_{c,j}$ from $\mathbf{V}^{s}_{b-1,c}$ ;
    // Across frames
     Compute semantic sensitivity:
    $\mathcal{I}(\mu_{c,j})=\frac{1}{\sigma^{2}_{c,j}}$ ;
    // Fisher Information
     Compute classification contribution:
    $\mathbb{E}[\gamma_{c,j}]\propto\frac{T_{c,j}\cdot\mu_{c,j}}{\lambda}$ ;
    // Cosine-aligned score
     Compute importance score:
    $I_{b-1,c,j}=\frac{T_{c,j}\cdot\mu_{c,j}}{\sigma^{2}_{c,j}}$ ;
    // Weighted relevance
Compute weighted distillation loss:
$\mathcal{L}_{FSSD}=\frac{1}{|\mathcal{D}^{b}|\cdot d_{vt}}\sum_{c}^{|\mathcal{C}_{b}|}\sum_{i}^{N_{c}}\sum_{j}^{d_{vt}}I_{b-1,c,j}\cdot\|\bar{V}_{b-1,c,i,j}^{s}-\bar{V}_{b,c,i,j}^{s}\|_{2}^{2}$
return $\mathcal{L}_{\text{FSSD}}$
Algorithm 1 Frame-Shared Semantics Distillation (FSSD)
Input: Video frames $\{\mathbf{x}^{\text{V}}_{i}\}_{i=1}^{N_{f}}$;
  Task-specific experts $\{\mathcal{G}_{k}\}_{k=1}^{K}$;
  Temporal anchors $\{\bar{\mathbf{V}}^{tem}_{c}\}_{c=1}^{C}$ for current task
Output: Final representation $\mathbf{V}$
3pt
Compute spatial mean: $\bar{\mathbf{V}}^{s}=\frac{1}{N_{f}}\sum_{i=1}^{N_{f}}\mathbf{x}^{\text{V}}_{i}$ ;
// Mean of frame features
for each expert $k=1$ to $K$ do
   Concatenate CLS token and frame features
 $\mathbf{V}_{k}^{st}=\mathcal{G}_{k}([\mathbf{x}^{\text{V}}_{\text{CLS}};\mathbf{x}^{\text{V}}_{1};\dots;\mathbf{x}^{\text{V}}_{N_{f}}])[0]$ ;
 // Spatiotemporal feature
  Compute temporal representation:
 $\mathbf{V}_{k}^{tem}=\mathbf{V}_{k}^{st}-\bar{\mathbf{V}}^{s}$ ;
 // Temporal decomposition
  Compute routing score:
 $r_{k}=\max_{c\in\mathcal{Y}_{k}}\cos(\mathbf{V}_{k}^{tem},\bar{\mathbf{V}}^{tem}_{c})$ ;
 // Similarity to temporal anchors
Compute final representation:
$\mathbf{V}=\bar{\mathbf{V}}^{s}+\sum_{k=1}^{K}r_{k}\cdot\mathbf{V}_{k}^{st}$ ;
// Residual fusion
return $\mathbf{V}$
Algorithm 2 Temporal Decomposition based Mixture-of-Experts Inference
Frame-Shared Semantics Distillation (FSSD).
Algorithm 1 mitigates forgetting by selectively preserving spatial feature channels that are semantically important and temporally stable across frames. Importance is computed per channel using two criteria: (1) Semantic sensitivity, quantified by Fisher Information, and (2) Classification contribution, measured by cosine similarity with text features. These weights are used in a weighted distillation loss between the frozen previous model and the current task model, enabling exemplar-free knowledge retention while allowing plasticity.
Temporal Decomposition-based Mixture-of-Experts (TD-MoE).
Algorithm 2 routes video inputs to task-specific experts based on temporal relevance. Each expert encodes spatiotemporal features from video frames, from which temporal dynamics are isolated via residual decomposition. Temporal features are then compared to precomputed class anchors to compute routing scores. Final representations are generated by combining the expert outputs with the spatial feature via residual fusion. This enables dynamic, task ID-agnostic inference driven by temporal structure.
Appendix B Theoretical Supplement
Semantic Sensitivity. The Fisher Information with respect to the mean parameter $\mu_{j}$ is defined as:
$\mathcal{I}_{j}(\mu_{c,j})=\mathbb{E}_{\bar{V}^{s}_{c,j}}\left[\left(\frac{\partial}{\partial\mu_{c,j}}\log p(\bar{V}^{s}_{c,j};\mu_{c,j})\right)^{2}\right],$ (16)
where $p(\bar{V}^{s}_{c,j};\mu_{c,j})$ is the probability density function of the Gaussian:
$p(\bar{V}^{s}_{c,j};\mu_{c,j})=\frac{1}{\sqrt{2\pi\sigma_{c,j}^{2}}}\exp\left(-\frac{(\bar{V}^{s}_{c,j}-\mu_{c,j})^{2}}{2\sigma_{c,j}^{2}}\right).$ (17)
Taking the derivative of the log-likelihood with respect to $\mu_{i}$:
$\log p(\bar{V}^{s}_{c,j};\mu_{c,j})=-\frac{1}{2}\log(2\pi\sigma_{c,j}^{2})-\frac{(\bar{V}^{s}_{c,j}-\mu_{c,j})^{2}}{2\sigma_{i}^{2}},$ (18)
$\frac{\partial}{\partial\mu_{c,j}}\log p(\bar{V}^{s}_{c,j};\mu_{c,j})=\frac{\bar{V}^{s}_{c,j}-\mu_{c,j}}{\sigma_{c,j}^{2}}.$ (19)
Then, the Fisher Information becomes:
$\mathcal{I}(\mu_{c,j})=\mathbb{E}\left[\left(\frac{\bar{V}^{s}_{c,j}-\mu_{c,j}}{\sigma_{c.j}^{2}}\right)^{2}\right]=\frac{1}{\sigma_{j}^{4}}\mathbb{E}\left[(\bar{V}^{s}_{c,j}-\mu_{c,j})^{2}\right]=\frac{1}{\sigma_{j}^{4}}\cdot\sigma_{c,j}^{2}=\frac{1}{\sigma_{c,j}^{2}}.$ (20)
Thus, we obtain:
$\mathcal{I}(\mu_{c,j})=\frac{1}{\sigma_{c,j}^{2}}.$ (21)
Classification Contribution. As we use cosine distance as classification basis, the $c$-th classification decision score $\gamma_{c,j}$ of the $j$-th channel is modeled as:
$\gamma_{c,j}=\frac{\bar{V}^{s}_{c,j}\cdot T_{c,j}}{\|\bar{\bm{V}}_{c}^{s}\|\cdot\|\bm{T}_{c}\|}\propto\frac{\bar{V}^{s}_{c,j}\cdot T_{c,j}}{\|\bar{\bm{V}}_{c}^{s}\|}.$ (22)
Therefore, the score function s can be approximately written as:
$\gamma_{c,j}=\sum_{j}\bar{V}^{s}_{c,j}\cdot\alpha_{j},\quad\alpha_{j}=\frac{T_{c,j}}{\|\bar{\bm{V}}_{c}^{s}\|}\approx\frac{T_{c,j}}{\lambda}.$ (23)
For the expected score $\mathbb{E}[\gamma_{c,j}]$, it can be calculated as:
$\mathbb{E}[\gamma_{c,j}]\propto\mathbb{E}\left[\frac{\sum_{i}\bar{V}^{s}_{c,j}\cdot T_{c,j}}{\|\bar{\mathbf{x}}_{v}\|}\right]=\sum_{j}{\alpha_{j}\cdot\mu_{i}}\approx\frac{T_{c,j}\cdot\mu_{c,j}}{\lambda}$ (24)
Then, the joint measure of informativeness is:
$I_{c,j}\propto\alpha_{j}\cdot\mu_{c,j}\cdot\mathcal{I}(\mu_{c,j})\approx\frac{T_{c,j}}{\lambda}\cdot\frac{\mu_{c,j}}{\sigma_{c,j}^{2}}.$ (25)
This expression provides a theoretically principled and interpretable metric for frame-shared semantics. It reflects the intuition that an informative channel should (i) be strongly activated on average ($\mu_{c,j}$ large), and (ii) exhibit consistent activation patterns across samples ($\sigma_{c,j}^{2}$ small). Therefore, we define the importance of channel $j$ as:
$I_{c,j}=\frac{T_{c,j}\cdot\mu_{c,j}}{\sigma_{c,j}^{2}}.$ (26)
This formulation also aligns with the the signal-to-noise ratio theory (SNR), providing a unified theoretical justification.
Appendix C Additional Experiments
C.1 Analysis of hyper-parameter.
Table 7: Analysis of hyperparameter $w$ on HMDB51 and UCF101 datasets. Best results are in bold.
Dataset Metric $\times 10^{3} 1×1$$\times 10^{4}$ $2.5\times 10^{4}$ $5\times 10^{4}$ $1\times 10^{5}$
HMDB($5\times 5$s) Acc $\uparrow$ 59.51 63.04 62.10 62.29 63.05
BWF $\downarrow$ 16.90 11.04 11.14 12.32 12.77
UCF101($5\times 10$s) Acc $\uparrow$ 85.10 85.79 84.54 84.88 84.16
BWF $\downarrow$ 7.93 5.63 6.95 5.67 6.83
Table. 7 analyzes the sensitivity of the hyper-parameter $w$, which controls the strength of FSSD regularization. Across a wide range of values, our framework maintains stable performance, indicating robustness to hyper-parameter variations. Moderate $w$ values achieve the best trade-off between knowledge retention and adaptability, avoiding under-regularization or excessive constraint on model plasticity.
C.2 Comparison of MoE Routing Strategies (task by task)
Figure 4: MoE Methods Performance per Task on UCF101 and HMDB51
[…]
We observe that TD-MoE consistently outperforms all baselines across datasets and task granularities. Its performance advantage becomes more pronounced as the number of tasks increases, reaching up to 6–8% improvement on later tasks. In contrast, Adapter-MoE and CLIP-MoE exhibit only marginal gains, which tend to saturate early, suggesting limited ability to model task-specific dynamics. These findings confirm that temporal decomposition is effective for guiding expert selection in a task ID-agnostic manner, and helps mitigate forgetting by capturing relevant spatiotemporal cues.
Figure 5: Heatmap of task selection by temporal decomposition-based router on the UCF101.
As shown in 5, temporal decomposition-based router is effective at task boundary decisions.
As shown in Figure 7, it maintains highest accuracy over time and effectively mitigates forgetting, demonstrating the complementary strengths of spatial semantic preservation and temporal dynamic modeling
C.3 Effectiveness of Frame-Shared Semantics Distillation (task by task)
Figure 6: Distillation Methods Performance per Task on UCF101 and HMDB51
Figure 6 compares the impact of different distillation strategies, including w/o Distillation, unified Distillation, and our proposed Frame-Shared Semantics Distillation (FSSD), on UCF101 and HMDB51. The vertical axis reports the relative improvement over the non-distillation baseline.
The results demonstrate that FSSD delivers the most consistent and significant improvements, particularly on HMDB51. While unified distillation offers slight improvements, it lacks consistency across tasks. This suggests that uniform constraints fail to address the heterogeneous semantic importance of feature channels. Moreover, the increasing gap between FSSD and other methods over time confirms that adaptive regularization based on frame-shared semantic importance is critical for preserving relevant knowledge across tasks in VCIL.
C.4 Task-by-task Ablation Study Visualization
Figure 7: Task-wise ablation analysis across incremental tasks on UCF101 and HMDB51.
As shown in Figure 7, we progressively add our modules (FSSD and TD-MOE) on the UCF101 and HMDB51 datasets. Significant improvements are observed at each incremental stage, with more pronounced gains as the number of tasks increases, especially in the long-term scenario (10 tasks). This further validates the effectiveness of our proposed method and the superior performance of our model.